### v1

# GMAI-VL & GMAI-VL-5.5M: A Large Vision-Language Model and A Comprehensive Multimodal Dataset Towards General Medical AI

Abstract
Despite significant advancements in general artificial intelligence, such as GPT-4, their effectiveness in the medical domain (general medical AI, GMAI) remains constrained due to the absence of specialized medical knowledge. To address this challenge, we present GMAI-VL-5.5M, a comprehensive multimodal medical dataset created by converting hundreds of specialized medical datasets into meticulously constructed image-text pairs. This dataset features comprehensive task coverage, diverse modalities, and high-quality image-text data. Building upon this multimodal dataset, we propose GMAI-VL, a general medical vision-language model with a progressively three-stage training strategy. This approach significantly enhances the model’s ability by integrating visual and textual information, thereby improving its ability to process multimodal data and support accurate diagnosis and clinical decision-making. Experimental evaluations demonstrate that GMAI-VL achieves state-of-the-art results across a wide range of multimodal medical tasks, such as visual question answering and medical image diagnosis. Our contributions include the development of the GMAI-VL-5.5M dataset, the introduction of the GMAI-VL model, and the establishment of new benchmarks in multiple medical domains. Code and dataset will be released at https://github.com/uni-medical/GMAI-VL.
1 Introduction
Large-scale Vision-Language Models (LVLMs) have rapidly evolved in recent years, effectively integrating visual perception with language understanding by leveraging large-scale multimodal data, which enables them to capture complex visual and textual patterns and drive significant advancements in image recognition, natural language processing, and multimodal tasks. With the advancement of multimodal integration technologies, the demand for high-precision processing of diverse data types in the medical field has become increasingly critical. The ability to effectively integrate and analyze various data modalities, such as medical images, clinical text, and structured clinical records, is pivotal for achieving accurate and comprehensive diagnostic and treatment decisions.
However, existing LVLMs, such as GPT-4 [1], are limited in medical applications due to their lack of domain-specific knowledge, highlighting the need for specialized solutions that effectively integrate medical expertise. Addressing this challenge requires constructing a comprehensive medical vision-language dataset and developing domain-specific models. For the medical dataset, it should provide high-quality medical knowledge, including the following three aspects:
Comprehensive medical tasks. To enhance the model’s applicability across various medical scenarios, the dataset should cover a wide range of medical contexts, such as disease types, symptoms, and treatments. Comprehensive task coverage can improve the model’s generalization ability and increase its reliability in real-world applications. However, existing models often focus on specific domains [28, 41, 72, 65, 35], limiting their broader applicability. Expanding the dataset’s scope will further enhance the model’s utility in clinical practice.
Rich multimodal representation. A well-rounded medical multimodal dataset should encompass various modalities, including different medical imaging types (such as CT, MRI, and X-rays) and diverse forms of textual data (such as medical records and imaging reports). This would allow models to better integrate multi-source information and improve their analytical capabilities. However, existing methods tend to focus on a single type of medical imagery [33, 70, 50], limiting the model’s adaptability to diverse clinical scenarios. A more diverse multimodal dataset would provide a foundation for developing more comprehensive models, better suited to the complexity of real-world medical environments.
Figure 1: Overview of GMAI-VL and GMAI-VL-5.5M. (a) illustrates the sources, departments, modalities, task types, and instruction formats of the GMAI-VL-5.5M dataset. (b) Architecture of GMAI-VL, integrating a Vision Encoder, Projector, and Large Language Model. (c) Three-stage training process of GMAI-VL, including shallow alignment, deep alignment, and instruction tuning with corresponding data sizes and training components. The flame symbol Refer to caption denotes the training part, while the snowflake symbol Refer to caption indicates frozen part.
High-quality image-text data. High-quality training data is essential for optimizing model performance, especially in medical applications. Ideally, image-text data for medical contexts should consist of a large, well-curated collection of medical images paired with precise textual descriptions. This data would enhance the model’s understanding of key medical concepts, such as diagnosis, treatment, and clinical workflows, ultimately contributing to improved clinical outcomes. While progress has been made with datasets sourced from platforms like PubMed [39, 54, 70], these datasets suffer from inconsistent data quality, imprecise alignment between images and text, and a lack of standardization, which limits their effectiveness.
Based on above observations, we propose an annotation-guided data generation methodology for developing a comprehensive multimodal medical datasets. The methodology begins by collecting large-scale, open-source medical imaging datasets and extracting key annotations, such as modality, task type, labels, and bounding boxes. A vision-language model (e.g., GPT-4o) is then used to transform these datasets, covering tasks like lesion detection, segmentation, and disease diagnosis, into high-quality image-text pairs for training LVLMs. To ensure data quality, extracted image information is incorporated into the prompt design, improving model performance across various clinical tasks. This results in a comprehensive multimodal dataset with 5.5M samples, named GMAI-VL-5.5M, which supports the development of general medical LVLMs. Fig. 1(a) illustrates the sources, departments, modalities, task types, and instruction formats of the constructed dataset.
With the constructed GMAI-VL-5.5M dataset, we develop a general medical vision-language model, GMAI-VL. To enhance its integration of visual and linguistic features and its instruction-following abilities, a three-stage training strategy is proposed in this paper. Specifically, we sequentially implement shallow and deep alignments in the first two stages, gradually building associations between visual (medical images) and language (medical texts) elements from basic features to high-level semantics. Next, we fine-tune the model with cross-modal instructions, improving its understanding of visual-language interactions and instruction-following in complex tasks. With this strategy, GMAI-VL shows strong performance in medical tasks like visual question answering and medical image diagnosis, providing a solid foundation for advancing multimodal models in the medical field.
The contributions of this work are as follows:
  • We build an annotation-guided data generation methodology to construct GMAI-VL-5.5M, a comprehensive vision-language dataset. This dataset offers extensive coverage of medical tasks, diverse multimodal representations, and high-quality image-text pairs, providing a robust foundation for model training.
  • Leveraging GMAI-VL-5.5M, we design GMAI-VL, a versatile medical vision-language model. Our proposed three-stage training strategy enhances its ability to integrate visual and linguistic features, significantly improving its performance in instruction-following and generalization across a wide range of medical tasks.
  • GMAI-VL outperforms existing models in multimodal question-answering tasks, including PMC-VQA and VQA-RAD. It sets new benchmarks on OmniMedVQA, GMAI-MMBench, and Health & Medicine track of MMMU. Specifically, GMAI-VL achieves an average score of 88.48% on OmniMedVQA, 62.43% on the GMAI-MMBench test set, and 51.3% on the Health & Medicine track of MMMU.
2 Related Work
Large-scale medical vision-language datasets of high quality and multiple modalities are the basis of Large Vision-Language Models (LVLMs) in the medical domain. While natural language and vision datasets are easily accessible online, biomedical datasets often focus on text or images only and many of them are limited to specific tasks or modalities, thus with unsatisfactory generalization ability. Notable datasets like MIMIC-CXR [33] and CheXpert [12] have advanced radiology models but are restricted to single image modality (X-ray), which hinders their use as general-purpose medical LVLMs. To address this, researchers have begun scraping public sources like PubMed and textbooks to construct large-scale vision-language datasets. Examples include datasets proposed in LLaVA-Med [39], Med-Flamingo [54], and PubMedVision [14], with PubMedVision optimizing LLaVA-Med dataset for higher-quality medical data. In addition to scraping, open-source image datasets with annotations can also be converted into image-text pairs for model training. Specifically, image information like modalities and annotations are input into large language models, e.g., GPT series, to generate text paired with the corresponding image. Some popular examples include the datasets constructed in RadFM [70], MedDr [28], MedTrinity-25M [71], ChiMed-VL [48], BiomedGPT [79], Med-Gemini [58], and Med-PaLM [59].
These efforts usually suffer from either limited modalities, data sources, or task coverage. Thus, their dataset quality needs further improvement. To this end, we propose to construct a comprehensive medical vision-language dataset with extensive coverage of medical tasks, diverse multimodal representations, and high-quality image-text pairs, forming a robust foundation for model training.
Medical vision-language models
are usually based on general-purpose Large Vision-Language Models (LVLMs). Most of them adapt LVLMs to specific medical applications using specialized medical datasets. For instance, Med-Flamingo [54] enhances OpenFlamingo-9B using 0.8 million interleaved and 1.6 million paired medical image-text data, highlighting the critical need for multimodal data in medical image analysis and automated report generation tasks. RadFM [70] improves PMC-LLaMA [69] by leveraging 16 million radiology images with text descriptions from diverse sources. Similarly, Med-PaLM [68] adapts PaLM-E [25] to the medical domain with approximately one million medical data samples, achieving state-of-the-art performance in diagnostic support and medical knowledge Q&A. LLaVA-Med [39] utilizes a large-scale biomedical figure-caption dataset extracted from PubMed Central to enhance LLaVA [66, 67] to better understand biomedical images and facilitate open-ended conversational interactions. Med-Gemini [58] leverages long-format question-answering datasets to improve the multimodal and long-contextual capabilities of the baseline Gemini model, enabling superior performance in complex medical Q&A and multimodal reasoning tasks. Additionally, HuatuoGPT-Vision [14] and MedDr [28] build medical multimodal datasets to adapt general-purpose LVLMs like LLaVA and InternVL to various medical modalities, including radiology, pathology, dermatology, and endoscopy.
Previous studies usually focus on constructing medical datasets to adapt general-purpose LVLMs but pay less attention to the adaptation strategies. However, naive training/adaptation strategies may not successfully adapt general-purpose LVLMs to the medical data, due to the large gap between the natural image-text pairs and the medical ones. Moreover, these strategies can hardly align the broad imaging modalities and various types of medical text (e.g., prescriptions, radiology reports, and electronic health records) to obtain generalizable features, thus limiting the models’ performance. Our work thus proposes a novel three-stage training strategy to better integrate the visual and language features to enhance generalization ability.
3 GMAI-VL-5.5M: A Comprehensive Multimodal Dataset
With rapid advancements in medical vision-language models (VLMs), constructing high-quality datasets is crucial for developing general-purpose models. Unlike previous approaches that primarily rely on published literature, our method utilizes specialized medical datasets to create a more robust, high-quality resource. We build GMAI-VL-5.5M, a comprehensive medical vision-language dataset that aggregates data from diverse open-source and proprietary sources. Covering 13 medical imaging modalities and 18 specialties, it addresses a wide range of medical imaging tasks. This dataset enhances the model’s ability to process complex medical information, advancing precision medicine and intelligent diagnostics.
Table 1: Comparison of various medical multimodal datasets, including details on the dataset size, modality type, language, data traceability, and sources of information.
Datasets Data Size Modality Language Traceability Data Source
PathVQA [29] 32.7k Pathology EN × Textbooks
MIMIC-CXR [33] 227k X-Ray EN Hospital
quilt-1M [32] 1M Pathology EN × YouTube & PubMed
MedDr VQA [28] 197k Multimodal EN 13 medical datasets
PMC-OA [43] 1.65M Multimodal EN × PubMed
PMC-VQA [80] 413k Multimodal EN × PubMed
LLaVA-Med VQA [39] 56,702 Multimodal EN × PubMed
ChiMed-VL [48] 1.05M Multimodal CN × PubMed
PMC-CaseReport [70] 438k Multimodal EN × PubMed
PubMedVision [14] 1.29M Multimodal EN&CN × PubMed
GMAI-VL-5.5M (ours) 5.5M Multimodal EN&CN 219 specialized medical imaging datasets
3.1 Data Curation
Data collection.
To construct a comprehensive multimodal medical dataset, we sourced 219 datasets from diverse platforms. Fig. 1(a) highlights key data sources, such as Kaggle, Grand Challenge, and Huggingface, which facilitate extensive data collection. These datasets cover diverse imaging modalities, including fundus, CT, MRI, and ultrasound (US), and span a variety of medical tasks, such as diagnosis, severity assessment, and organ recognition. They also encompass multiple clinical specialties, including pathology, dermatology, ophthalmology, otolaryngology, and oncology, further enhancing their diversity.
Data preprocess.
After data collection, we apply a preprocessing workflow to extract 2D medical images from the videos and 3D medical volumes. For data cleaning, we first extract key annotation information from classification data, including modality, department, and labels for each image, discarding instances with missing or unclear annotations. For segmentation data, we follow the SA-Med2D-20M [75] preprocessing method, filtering out low-quality images and labels, and converting them into detection datasets. The converted segmentation data is then processed together with the original detection dataset, with key annotation information extracted and linked to each image. Finally, the preprocessed data are then standardized and organized into a structured format: <image, modality, label, department, bbox [optional]>.
Data generation.
Subsequently, the data are categorized into two primary types: classification datasets and detection/segmentation datasets. Each category is further refined using specific prompts tailored for large model training. For data generation, large vision-language models (GPT-4o) are employed to produce detailed image descriptions and corresponding instruction-following data through annotation-guided data generation methodology. For classification datasets, detailed descriptions of the entire image are generated, while for detection datasets, the focus is on specific regions enclosed by bounding boxes, providing comprehensive functional analyses of these areas. Notably, the segmentation dataset was transformed into a detection dataset using external bounding boxes, and data generation followed detection dataset protocols. Furthermore, to improve the model’s multilingual capability, we translated a portion of English image-text data into Chinese. Incorporating multilingual data helps to enhance the generalization capabilities of domain-specific multimodal models. The resulted data is utilized for medical Visual Question Answering (VQA) tasks, forming the comprehensive VQA dataset, named GMAI-VL-5.5M. The detailed pipeline for generating prompt-driven data is illustrated in Fig. 2.
Figure 2: The prompt-driven data generation pipeline comparing without-annotation-guided and annotation-guided methods. The annotation-guided approach integrates specific annotation information (e.g., <image, modality, label, department, bbox [optional]>) to generate high-quality, accurate descriptions, while the without-annotation-guided approach often results in lower-quality outputs. Figure with complete prompt and response is provided in Supp. Mat..
As depicted in Fig. 1(a), it contains six kinds of instrcuction-following formats, including image-level captions, region-level captions, free instructions, dialogue, visual perception and text-only tasks. The specific composition of GMAI-VL-5.5M can be found in Supp. Mat. (Table 6). These formats enable VLMs to better understand and process complex visual and textual information in medical contexts. The GMAI-VL-5.5M dataset significantly enhances the model’s cross-modal reasoning ability, enabling it to handle complex multimodal inputs in real clinical scenarios. The richness of the instruction formats allow the model to progress from basic question answering to advanced medical image analysis, ultimately providing strong support for clinical diagnosis and decision-making.
3.2 Data Property
Data statistics.
These datasets encompass diverse medical imaging tasks and modalities, forming a solid foundation for developing and evaluating medical LVLMs. Fig. 4 (in Supp. Mat.) illustrates the distribution of modalities, tasks, clinical departments, and specific medical challenges represented within the collected datasets. This visualization highlights the extensive diversity and coverage of our data collection efforts. After careful standardization and integration, these datasets form the core of our comprehensive medical image-text dataset, GMAI-VL-5.5M, which serves as a crucial resource for advancing precision medicine and intelligent diagnostic systems.
Accuracy and reliability of generated data.
The accuracy and reliability of the generated data are primarily addressed from two key aspects. First, the reliability of the data source is ensured by utilizing high-quality datasets from reputable sources. The data used for generation primarily comes from professional medical challenges (such as Kaggle and GrandChallenge) and publicly available, peer-reviewed datasets. This guarantees the integrity and accuracy of the data at its source. Second, the data generation process is rigorously controlled to ensure accuracy and reliability. Although GPT was used for data generation, prompts were meticulously crafted to include essential annotations, minimizing errors and hallucinations. As illustrated in Fig. 2, the proposed annotation-guided approach incorporates specific annotations (e.g., <image, modality, label, department, bbox [optional]>) to produce precise, high-quality descriptions. Compared to the without-annotation-guided approach, which often results in less detailed outputs, the annotation-guided method delivers more professional, pathology-specific descriptions, significantly enhancing the reliability of the generated data.
Compared with other medical multimodal datasets.
The GMAI-VL-5.5M dataset, as highlighted in Table 1, stands out due to its unmatched scale, encompassing over 5.5 million samples from more than 219 specialized medical imaging datasets. Unlike other datasets listed, GMAI-VL-5.5M supports a wider variety of modalities and languages, making it a truly global resource that caters to diverse clinical needs. Additionally, GMAI-VL-5.5M emphasizes traceability of its data, ensuring a high standard of clinical relevance and reliability. This comprehensive and diverse dataset is critical for pushing the boundaries of medical multimodal research, enabling more effective training of LVLMs that can generalize across multiple medical tasks and scenarios, thereby driving innovations in precision medicine and intelligent diagnostics.
4 GMAI-VL: A General Medical Vision-Language Model
4.1 Architecture
The GMAI-VL model is a vision-language model built upon the LLaVA architecture [45, 39], incorporating three key components: a large language model (LLM), a vision encoder, and a projector (MLP), as illustrated in Fig. 1(b). These components are designed to work together seamlessly, enabling the model to deliver exceptional performance in medical applications.
We adopt InternLM2.5-7B [63] as our language processing module, which offers outstanding reasoning capabilities. With a context window up to one million tokens, it can handle complex medical tasks and generate coherent, accurate responses. Its support for advanced instruction-following makes it particularly effective in addressing intricate medical queries, thereby enhancing the model’s ability to understand and respond to a wide range of instructions.
For vision processing, GMAI-VL employs a CLIP-based vision encoder [55], which transforms visual inputs into high-dimensional feature representations. CLIP’s strong performance in aligning image and text representations ensures that medical image features are accurately extracted and effectively integrated with linguistic information, significantly enhancing the model’s ability to handle multimodal medical data.
The MLP, as a projector, serves as a bridge between the vision encoder and the LLM, optimizing high-dimensional outputs and further enhancing feature representation. The seamless integration of these components enables GMAI-VL to excel in processing and understanding complex medical multimodal data.
4.2 Optimization Strategy
As illustrated in Fig. 1(c), the training process of the GMAI-VL model is divided into three stages: shadow alignment, deep alignment, instruction tuning, respectively. The detailed hyper-parameter settings can be found in Supp. Mat. (Table 8). To enhance the training of GMAI-VL, we supplement our GMAI-VL-5.5M dataset with additional medical datasets, ultimately forming an 11.7M dataset for training. This supplemented data increases the diversity of training data, exposing the model to a wider range of medical scenarios and visual-language patterns, enhancing its generalization to complex clinical tasks and ensuring robustness in real-world applications. Fig. 5 (in Supp. Mat.) provides a complete distribution of the utilized datasets during the training stage. The detailed data proportions for each training stage are detailed in Table 7 (in Supp. Mat.).
Stage I: Shadow alignment.
In the shallow alignment phase, we utilize a large-scale medical image-text dataset comprising approximately 11.7 million image-text pairs, sourced from a combination of publicly accessible datasets and proprietary in-house data. To achieve shallow alignment, we freeze both the large language model and the vision encoder, optimizing only the projector. With this optimization stage, the model establishes an initial alignment between medical images and their corresponding textual descriptions. All input images are resized to $336\times 336$ pixels, and the training objective is to minimize the cross-entropy loss of the text tokens.
Stage II: Deep alignment.
Most vision encoders in multimodal models are pre-trained on natural images, so it is essential to address the domain gap between medical and natural images during the deep alignment stage. To achieve this, we fine-tune both the vision-language projector and the vision encoder, enhancing the alignment between the visual features of medical images and the language model’s feature space.
Stage III: Instruction tuning.
At this stage, we fine-tune our GMAI-VL model (including the vision encoder, the language model, and the projector components) by instruction tuning to enhance its instruction-following and dialogue capabilities. The multimodal instruction tuning data is primarily derived from the training data in previous stages, with a filtering process to select high-quality and more suitable data for fine-tuning. Specifically, we manually reviewed a subset of each dataset and labeled it as either “high quality” or “low quality.” High-quality datasets were used for tuning, while most low-quality ones were excluded, except for a small portion retained for diversity. This approach allowed efficient filtering without reviewing every sample. Additionally, we incorporate medical text dialogue data to ensure the model’s versatility in handling various dialogue scenarios. Thus, our instruction tuning data comprises approximately ten million samples.
5 Experimental Results
To evaluate our model, we employed several established multimodal medical benchmarks, each targeting specific aspects of medical image understanding and question answering. Below is a brief overview of the benchmarks used in our experiments:
  • Traditional medical VQA benchmarks: Traditional multimodal medical question-answering benchmarks, such as VQA-RAD [38], SLAKE [44], and PMCVQA [80], span various imaging modalities tasks, primarily assessing the model’s ability to extract information from medical images and answer clinical questions. They evaluate the model’s performance in understanding medical imaging and integrating multimodal information.
  • OmniMedVQA: OmniMedVQA [31] provides a rich dataset of paired medical images and text, designed to evaluate the model’s ability to recognize and understand fundamental medical imaging concepts, focusing on cross-modal reasoning and information integration.
  • GMAI-MMBench: GMAI-MMBench [16] focuses on assessing the model’s ability to identify fine-grained objects in complex clinical scenarios, challenging its capacity to handle long-context tasks and accurately recognize and reason over detailed medical features.
  • MMMU Health & Medicine track: The Health & Medicine track of the MMMU [78] benchmark spans a wide range of medical fields, derived from university exams, quizzes, and textbooks. It evaluates the model’s reasoning ability in complex medical scenarios and the specialized knowledge in health and medicine.
5.1 Medical VQA Benchmarks
Table 2: Results on Traditional Medical VQA Benchmarks. The highest performance in each column is highlighted in red, and the second-highest performance is highlighted in blue.
Model VQA-RAD SLAKE PMC-VQA Avg.
Med-Flamingo [54] 45.4 43.5 23.3 37.4
RadFM [70] 50.6 34.6 25.9 37.0
LLAVA-Med-7B [39] 51.4 48.6 24.7 41.6
Qwen-VL-Chat [6] 47.0 56.0 36.6 46.5
Yi-VL-34B [77] 53.0 58.9 39.5 50.5
LLAVA-v1.6-7B [46] 52.6 57.9 35.5 48.7
LLAVA-v1.6-13B [46] 55.8 58.9 36.6 50.8
LLAVA-v1.6-34B [46] 58.6 67.3 44.4 56.8
HuatuoGPT-Vision-7B [14] 63.8 74.5 52.7 63.7
GMAI-VL(w/o our data) 62.3 66.3 39.0 55.9
GMAI-VL(ours) 66.3 72.9 54.3 64.5
The performance of various VLMs on popular medical VQA benchmark datasets is summarized in Table 2, including VQA-RAD[38], SLAKE[44], and PMC-VQA[80]. Our model, GMAI-VL, demonstrates a strong performance, achieving the highest score on the VQA-RAD[38] dataset with 66.3%, outperforming other models such as HuatuoGPT-Vision-7B. This result highlights GMAI-VL’s superior capability in handling radiological image question-answering tasks. For the PMC-VQA[80] dataset, GMAI-VL achieves 54.3%, and 72.9% on SLAKE[44], demonstrating its capability in handling medical VQA tasks across diverse modalities.
GMAI-VL demonstrates competitive performance across multiple benchmarks, showcasing its versatility in medical image understanding and question-answering.
5.2 OmniMedVQA
Table 3: Comparison of performance between representative LVLMs and GMAI-VL on OmniMedVQA across five different question type. The best performance in each column is highlighted in red, and the second-best performance is highlighted in blue. Abbreviations: MR = Modality Recognition, AI = Anatomy Identification, DD = Disease Diagnosis, LG = Lesion Grading, OBA = Other Biological Attributes.
Model MR AI DD LG OBA Overall
Random Guess 25.00 25.84 28.41 25.40 37.49 28.28
Open-Source LVLMs
MiniGPT-4 [81] 36.98 32.68 24.19 20.45 26.14 27.59
LLaVA [45] 52.30 35.27 11.80 9.77 24.70 22.86
LLaMA_Adapter_v2 [27] 58.45 38.18 29.12 23.73 30.97 35.08
InstructBLIP [20] 72.35 39.90 32.01 43.80 47.91 41.14
BLIP-2 [40] 57.48 49.83 46.21 30.52 73.52 50.77
Qwen-VL-Chat [6] 33.69 10.95 16.27 6.71 41.68 20.29
mPLUG-Owl2 [76] 78.01 48.52 39.68 20.56 59.36 48.44
LLaVa-NeXT [46] 68.23 46.74 41.21 18.43 39.57 45.57
DeepSeek-VL [49] 74.01 51.94 45.46 21.06 29.04 48.76
Yi-VL [77] 59.56 44.81 48.97 32.93 24.63 47.28
InternVL2-40B [18] 96.76 64.25 76.28 76.50 76.27 78.70
Medical Special Model
MedVInT-TE [80] 62.62 41.03 40.57 12.17 45.17 43.83
LLaVA-Med [39] 48.41 27.96 23.72 16.10 21.94 27.82
Med-Flamingo [54] 26.74 25.10 23.80 28.04 16.26 23.82
RadFM [70] 27.45 21.65 23.75 16.94 20.05 23.48
MedDr [28] 91.37 51.62 65.56 73.18 74.52 68.27
HuatuoGPT-Vision-34B [14] 95.06 75.67 66.51 72.83 74.92 73.23
Our Model
GMAI-VL(w/o our data) 96.40 80.97 79.14 70.29 75.66 79.96
GMAI-VL(ours) 98.64 92.95 88.7 87.21 82.95 88.48
Table 3 summarizes the performance of various large vision-language models (LVLMs), including our proposed GMAI-VL, across five question types: Modality Recognition, Anatomy Identification, Disease Diagnosis, Lesion Grading, and Other Biological Attributes. GMAI-VL demonstrates outstanding accuracy across multiple tasks, achieving 98.64% in Modality Recognition, 92.95% in Anatomy Identification, and 88.71% in Disease Diagnosis. It outperforms both open-source LVLMs and medical-specific models, underscoring its capability to accurately identify anatomical structures and diagnose diseases from visual data. In Lesion Grading, GMAI-VL attained the highest score of 87.21%, and it also delivers a strong performance of 82.95% in Other Biological Attributes, showcasing its versatility across diverse biological contexts. With an average accuracy of 88.48%, the highest among all evaluated models, GMAI-VL excels not only in general medical question-answer tasks but also in complex reasoning requiring domain-specific knowledge, surpassing models like HuatuoGPT-Vision-34B and InternVL2-40B.
These results verify our GMAI-VL is a leading model in multimodal medical image understanding, setting a new benchmark for medical VQA tasks. Its consistent top performance across question types highlights its potential for broader applications in medical question answering.
5.3 GMAI-MMBench
Table 4: Results on the val and test sets of GMAI-MMBench for clinical VQA tasks. The full names of the evaluated tasks can be found in Table 5 in literature [16]. The best model in each category is highlighted in red, while the second-best model is indicated in blue.
Model Name Overall (val) Overall (test) AR BVR B CR C DD IQG MR M NT OR-A OR-HN OR-P OR-T SG SAR SIR SWR
Random Guess
Random 25.70 25.94 38.20 22.73 22.92 22.72 24.06 26.66 27.13 27.00 20.00 24.75 21.37 22.93 22.33 21.18 32.43 24.23 21.39 23.71
Open-Source LVLMs
Flamingo v2 [4] 25.58 26.34 37.74 21.50 20.62 22.00 22.41 27.29 25.91 27.45 18.00 28.79 25.16 22.13 22.00 22.00 34.61 22.88 20.44 27.43
VisualGLM-6B [22] 29.58 30.45 40.16 33.92 24.92 25.22 24.21 32.99 29.96 29.53 21.20 37.88 30.32 24.80 13.33 29.88 33.11 19.62 19.16 37.43
InstructBLIP-7B  [20] 31.80 30.95 42.12 26.92 24.92 28.09 21.65 34.58 31.58 29.23 22.40 30.30 28.95 27.47 23.00 24.82 32.88 19.81 21.64 26.57
Qwen-VL  [6] 34.80 36.05 37.05 37.24 35.85 28.98 24.81 43.60 24.70 30.12 19.20 44.44 29.68 31.87 25.00 31.18 30.26 21.54 20.10 26.86
Yi-VL-6B  [77] 34.82 34.31 41.66 39.16 26.62 30.23 31.88 38.01 26.72 24.93 25.20 37.37 29.58 31.20 32.33 30.59 36.71 24.81 23.18 31.43
ShareGPT4V-7B  [15] 36.71 36.70 43.96 37.59 21.54 37.57 18.80 43.26 32.39 27.30 22.80 43.43 29.47 37.33 22.00 31.76 34.98 24.42 25.06 30.00
LLAVA-V1.5-7B [45] 38.23 37.96 45.45 34.27 30.92 41.32 21.65 44.68 34.01 27.74 23.60 43.43 28.00 42.13 29.00 35.06 33.41 22.12 23.61 29.14
XComposer2  [24] 38.68 39.20 41.89 37.59 33.69 40.79 22.26 45.87 36.44 32.94 27.20 58.59 26.11 36.40 43.67 37.29 32.06 23.46 27.80 32.86
LLAVA-InternLM-7b [19] 38.71 39.11 36.36 36.54 32.62 38.10 30.68 46.53 34.82 28.19 25.20 48.99 28.11 40.53 33.33 36.00 34.08 26.73 24.12 29.71
InternVL-Chat-V1.5 [18] 38.86 39.73 43.84 44.58 34.00 33.99 31.28 45.59 33.20 38.28 32.40 42.42 31.89 42.80 27.00 36.82 34.76 23.27 24.72 32.57
InternVL-Chat-V1.2  [17] 39.52 40.01 41.66 44.06 27.38 38.46 34.29 46.99 33.60 34.42 21.20 47.98 30.63 42.80 27.67 35.88 35.59 23.85 24.98 28.00
LLAVA-InternLM2-7b [19] 40.07 40.45 39.82 37.94 30.62 35.24 29.77 48.97 34.01 25.96 20.80 53.03 30.95 42.67 32.00 39.88 32.43 21.73 24.38 38.00
DeepSeek-VL-7B  [49] 41.73 43.43 38.43 47.03 42.31 37.03 26.47 51.11 33.20 31.16 26.00 44.95 36.00 58.13 36.33 47.29 34.91 18.08 25.49 39.43
MiniCPM-V2 [73] 41.79 42.54 40.74 43.01 36.46 37.57 27.82 51.08 28.74 29.08 26.80 47.47 37.05 46.40 25.33 46.59 35.89 22.31 23.44 31.71
Proprietary LVLMs
Claude3-Opus [2] 32.37 32.44 1.61 39.51 34.31 31.66 12.63 39.26 28.74 30.86 22.40 37.37 25.79 41.07 29.33 33.18 31.31 21.35 23.87 4.00
Qwen-VL-Max  [6] 41.34 42.16 32.68 44.58 31.38 40.79 10.68 50.53 32.79 44.36 29.20 51.52 41.37 58.00 30.67 41.65 26.95 25.00 24.64 39.14
GPT-4V  [1] 42.50 44.08 29.92 48.95 44.00 37.39 12.93 52.88 32.79 44.21 32.80 63.64 39.89 54.13 37.00 50.59 27.55 23.08 25.75 37.43
Gemini 1.0  [62] 44.38 44.93 42.12 45.10 46.46 37.57 20.45 53.29 35.22 36.94 25.20 51.01 34.74 59.60 34.00 50.00 36.64 23.65 23.87 35.43
Gemini 1.5  [56] 47.42 48.36 43.50 56.12 51.23 47.58 2.26 55.33 38.87 48.07 30.00 76.26 51.05 75.87 46.33 62.24 20.57 27.69 30.54 40.57
GPT-4o  [1] 53.53 53.96 38.32 61.01 57.08 49.02 46.62 61.45 46.56 56.38 34.00 75.25 53.79 69.47 48.67 65.88 33.93 22.88 29.51 39.43
Medical Special Model
Med-Flamingo  [54] 12.74 11.64 6.67 10.14 9.23 11.27 6.62 13.43 12.15 6.38 8.00 18.18 9.26 18.27 11.00 11.53 12.16 5.19 8.47 11.43
LLaVA-Med  [39] 20.54 19.60 24.51 17.83 17.08 19.86 15.04 19.81 20.24 21.51 13.20 15.15 20.42 23.73 17.67 19.65 21.70 19.81 14.11 20.86
Qilin-Med-VL-Chat  [48] 22.34 22.06 29.57 19.41 16.46 23.79 15.79 24.19 21.86 16.62 7.20 13.64 24.00 14.67 12.67 15.53 26.13 24.42 17.37 25.71
RadFM  [70] 22.95 22.93 27.16 20.63 13.23 19.14 20.45 24.51 23.48 22.85 15.60 16.16 14.32 24.93 17.33 21.53 29.73 17.12 19.59 31.14
MedDr [28] 41.95 43.69 41.20 50.70 37.85 29.87 28.27 52.53 36.03 31.45 29.60 47.47 33.37 51.33 32.67 44.47 35.14 25.19 25.58 32.29
Our Model
GMAI-VL(w/o our data) 54.99 56.23 51.26 61.05 53.79 44.39 44.51 62.60 40.80 57.42 35.20 79.50 61.31 77.81 53.60 69.29 35.39 35.77 29.71 44.86
GMAI-VL(ours) 61.74 62.43 75.26 59.66 67.24 56.86 54.29 67.14 42.80 79.97 41.60 75.00 60.45 75.48 53.33 58.12 42.09 72.31 37.40 59.14
The GMAI-MMBench benchmark is a comprehensive medical multimodal benchmark designed to evaluate models on a range of clinical visual question-answering (VQA) tasks. Table 4 presents the results of various LVLMs, including open-source LVLMs and commercial models, evaluated on the val and test sets across multiple clinical tasks.
Notably, GMAI-VL excels in specific tasks such as abnormality recognition (73.78%), biological variation recognition (63.06%), and clinical disease diagnosis (66.67%). These results demonstrates the model’s strong ability in understanding and interpreting complex clinical images. In comparison to other models, GMAI-VL consistently achieves either the best or second-best performance across most tasks. For instance, it ranks first in 16 out of 20 categories, including key tasks such as AR (Attribute Recognition) and DD (Disease Diagnosis), where it achieved scores of 75.26% and 67.14%, respectively, suggesting GMAI-VL’s strength in understanding medical scenarios.
Overall, GMAI-VL establishes a new benchmark in various clinical VQA tasks, demonstrating its potential as a reliable and versatile tool in medical multimodal applications.
5.4 MMMU Health & Medicine track
The MMMU benchmark, a widely recognized standard for evaluating multimodal models, was used to assess the performance of our proposed GMAI-VL model on the Health & Medicine track. The experimental results in Table 5 show the model’s performance across five key categories: Basic Medical Science (BMS), Clinical Medicine (CM), Diagnostics and Laboratory Medicine (DLM), Pharmacy (P), and Public Health (PH). GMAI-VL performs strongly across multiple categories, achieving top scores in DLM (43.3%), P (50.0%), and PH (53.3%), surpassing competitive models like LLaVA-v1.6 and HuatuoGPT-Vision-7B. These results highlight the model’s proficiency in handling complex tasks requiring diagnostic reasoning, pharmaceutical knowledge, and public health expertise. In BMS, GMAI-VL scores 50.0%, achieve the best performance, demonstrating the model’s the capacity of understanding medical knowledge. In CM, the model achieves 60.0%, remaining competitive with other leading models. These results underscore the model’s ability in processing both clinical and foundational medical information effectively. Overall, GMAI-VL achieves an average score of 51.3% across the Health & Medicine track, which is a top performance among other models, verifying its versatility in specialized medical domains.
Table 5: Performance on the val set for the MMMU Health & Medicine track. This track is divided into five categories: BMS (Basic Medical Science), CM (Clinical Medicine), DLM (Diagnostics and Laboratory Medicine), P (Pharmacy), and PH (Public Health). The best performance in each column is highlighted in red, and the second-best performance is highlighted in blue.
Model BMS CM DLM P PH MMMU Health & Medicine
Med-Flamingo [54] 33.6 30.2 23.3 29.3 25.8 28.4
RadFM [70] 31.6 28.6 26.7 26.2 26.8 27.9
LLaVA-Med-7B [39] 33.8 32.3 26.7 40.7 43.3 38.6
Qwen-VL-Chat [6] 32.7 20.6 19.3 29.6 33.3 31.7
Yi-VL-34B [77] 48.1 55.6 36.7 35.4 31.3 48.2
LLaVA-v1.6-7B [45] 46.4 43.4 30.0 29.6 26.7 33.1
LLaVA-v1.6-13B [45] 53.6 46.7 33.3 22.2 40.0 39.3
HuatouGPT-Vision-7B [14] 50.0 63.3 36.7 48.1 53.3 50.3
GMAI-VL(w/o our data) 43.3 56.7 43.3 46.7 40.0 46.0
GMAI-VL(ours) 50.0 60.0 43.3 50.0 53.3 51.3
5.5 Ablation Study on GMAI-VL-5.5M Dataset
We further analyzed the effectiveness of the proposed GMAI-VL-5.5M dataset. By removing GMAI-VL-5.5M from the training set, we observed a significant performance drop, as shown in Table 2, Table 3, Table 4 and Table 5. In the VQA-RAD, SLAKE, and PMC-VQA benchmarks, removing GMAI-VL resulted in an average performance improvement of 8.6%. For the OmniMedVQA, GMAI-MMBench, and MMMU Health & Medicine track, the improvements were 8.52%, 6.75%, and 5.3%, respectively. These results demonstrate that the GMAI-VL-5.5M dataset provides highly accurate and reliable medical knowledge, particularly in recognizing and understanding multimodal medical data, which significantly boosts model performance. This underscores the dataset’s diversity, comprehensiveness, and its ability to effectively complement other datasets in complex medical tasks.
6 Conclusion
In this study, we develop GMAI-VL, a large vision-language model, along with GMAI-VL-5.5M, a comprehensive multimodal medical dataset aimed at advancing general medical AI (GMAI). GMAI-VL-5.5M, which converts hundreds of medical image analysis datasets into high-quality image-text pairs through annotation-guided data generation, enables GMAI-VL to effectively address a wide range of clinical tasks. Experimental results show that GMAI-VL-5.5M significantly enhances GMAI-VL’s performance on diverse clinical tasks, achieving state-of-the-art results across several key benchmarks.
Supplementary Material
7 Expanded Data Generation Example
In this section, we present the full version of Fig. 2, which provides a detailed example of the data generation process, as shown in Fig.3. This expanded illustration includes the complete prompt and the corresponding detailed responses, enabling a clear comparison between data generation methodologies with and without annotation guidance. The detailed comparison allows for a more intuitive understanding of the distinctions between both methods.
Figure 3: The full version of Fig. 2 in the main text illustrates the complete of data generation pipeline, comparing without-annotation-guided and annotation-guided methods. The annotation-guided approach integrates specific annotation information (e.g., <image, modality, label, department, bbox [optional]>) to generate high-quality, accurate descriptions, while the without-annotation-guided approach often results in lower-quality outputs.
8 Details of GMAI-VL-5.5M
In Table. 6, we provide sub-datasets information of the multimodal dataset GMAI-VL-5.5M we have constructed. Based on the different data forms introduced in the paper, we have categorized the data into five distinct sub-datasets. These include GMAI-MM-Caption-1.7M, GMAI-MM-Instruct-0.9M, GMAI-MM-Percept-1.3M, GMAI-Text-Single-1M, and GMAI-Text-0.6M. Each sub-dataset corresponds to specific components: image caption data, free instruction data, visual perception data, text-only, and conversation. Furthermore, Fig. 4 presents a comprehensive distribution of the data across different modalities, original tasks, departments, and clinical task within this multimodal dataset, highlighting its extensive richness and diversity.
Table 6: Sub-Dataset Details for GMAI-VL-5.5M
Dataset Sub-Dataset Name Description Size
GMAI-VL-5.5M GMAI-MM-Caption-1.7M A curated set of detailed medical image captions. 1.7M
GMAI-MM-Instrunct-0.9M A diverse set of instructions for medical image analysis. 0.9M
GMAI-MM-Percept-1.3M A dataset of labels for medical image classification and segmentation. 1.3M
GMAI-Text-Single-1M A set of single-round medical dialogues on patient queries 1.0M
GMAI-Text-Multi-0.6M A dataset of multi-turn medical conversations on various topics. 0.6M
9 Training Data Overview
In this section, we provide a comprehensive overview of all datasets utilized for training the GMAI-VL model. The details include the dataset names, their corresponding categories, the amount of data used for training, and the proportion of training data allocated to each dataset during the three phases of model training. Fig. 5 visualizes the distribution of all our training data, showing the proportion of each category and subcategory.
Table. 7 summarizes the datasets employed, highlighting their respective categories and sizes. It is important to note that for certain datasets, we performed data cleaning and bilingual translation. As a result, the dataset sizes reported here may differ from the official numbers.
(a) (a) Modality distribution
(b) (b) Original task distribution
(c) (c) Department distribution
(d) (d) Clinical task distribution
Figure 4: Distribution of GMAI-VL-5.5M across tasks, modalities, departments, and clinical tasks. (a) Original Task Distribution: The dataset includes 2D Classification (50.4%), 3D Segmentation (30.3%), 2D Segmentation (12.7%), and 2D Detection (6.6%). (b) Modality Distribution: In addition to CT (26.8%) and MR (24.7%), X-ray (12.6%), Pathology (11.2%), and less common modalities like Dermoscopy (3.5%), Microscopy (2.4%), and PET (0.2%) are represented. (c) Department Distribution: While Orthopedic Surgery (12.9%) and General Surgery (10.3%) are the top contributors, departments like Endocrinology (1.3%), Infectious Diseases (0.8%), and Urology (0.7%) also provide data. (d) Clinical Task Distribution: Besides Disease Diagnosis (40.4%) and Organ Recognition (16.0%), tasks such as Muscle Recognition (3.3%), Nervous Tissue Recognition (1.5%), and Microorganism Recognition (1.2%) are included. Note: The distribution statistics shown here pertain only to the multimodal components of GMAI-VL-5.5M.
Figure 5: Distribution of all our training data. The inner ring represents major categories, each depicted in a different color. The outer ring corresponds to the subcategories within each major category. The size of each segment is proportional to the amount of data, as indicated in the legend, where the data volume for each subcategory is also provided.
Table 7: List of datasets used in our model. We employ a large collection of image-text data and instruction data for training stage.
Dataset Category Dataset Name Size ratio in stage 1&2 ratio in stage 3
General Captioning ALLaVA[13] 468k 100.0% 50.0%
ShareGPT4V[15] 102k
Medical Captioning GMAI-MM-Caption-1.7M 1.7M 100.0% 100.0%
PubMedVision[14] 1.3M
MedICaT[60] 173k 100.0% 5.0%
MPx-Single[70] 31k
PMC-OA[43] 1.3M
QUILT-1M[32] 643k
Retina Image Bank[3] 22k
Report Generation CheXpertPlus[12] 223k 100.0% 30.0%
MIMIC-CXR[33] 486k
OpenI[21] 7k
General Instruction GeoQA+[11] 72k 100.0% 75.0%
AI2D[36] 12k
SynthDoG[37] 29k
ChartQA[51] 18k
MMChemExam[42] 219k
LLaVA-Instruct-150K[45] 157k
DVQA[34] 200k
DocVQA[52] 10k
Medical Instruction GMAI-MM-Percept-1.3M 1.3M 100.0% 100.0%
GMAI-MM-Instruct-0.9M 0.9M
PubMedVision[14] 1.28M
LLaVA-Med-60k[39] 56k
PMC-Inline[70] 288k 100.0% 10.0%
VQA-Med-2019[8] 3.2k
Medical-Diff-VQA[30] 260k
PathVQA[29] 2.6k
PMC-CaseReport[70] 109k
PMC-VQA[80] 251k
ROCOV2[57] 60k
SLAKE[44] 0.6k
VQA-RAD[38] 0.3k
General Text blossom_orca[5] 20k 0.0% 100.0%
COIG-CQIA[7] 14.8k
Cosmopedia-100k[9] 33k
ShareGPT4V[15] 26k
Orca-Math[53] 379k
Leetcode[10] 1.7k
LogiQA[47] 12.7k
Lima[26] 83k
Open Hermes 2.5[64] 200k
Firefly[74] 189k
UltraChat[23] 189k
Alpaca-Instruct-52K[61] 49k
Medical Text GMAI-Text-Single-1M 1.0M 0.0% 100.0%
GMAI-Text-Multi-0.6M 649k
Overall - 15.7M - -
Figure 6: Diagram of the three-stage training process.
10 Model Training Settings
Table 8: Training settings of GMAI-VL’s stage I, stage II, and stage III.
Settings Stage I Stage II Stage III
freeze LLM True True False
freeze MLP False False False
freeze Vision Encoder True False False
packing type soft packing soft packing soft packing
learning rate 1e-3 1e-4 1e-5
learning rate schedule cosine decay cosine decay cosine decay
optimizer AdamW AdamW AdamW
optimizer hyper-parameters $\beta_{1}=0.9,\beta_{2}=0.999$ $\beta_{1}=0.9,\beta_{2}=0.999$ $\beta_{1}=0.9,\beta_{2}=0.999$
input size 336x336 336x336 336x336
total batch size 32x8x2 32x4x4 32x4x4
drop rate 0.0 0.0 0.0
numerical precision DeepSpeed bf16 DeepSpeed bf16 DeepSpeed bf16
GPUs for training 32xA100 (80G) 32xA100 (80G) 32xA100 (80G)
The table. 8 presents the training settings for GMAI-VL across three stages, detailing key hyperparameters. The train diagram as illustrated in Fig. 6.
  • Stage I (Shallow alignment). During this stage, both the large language model (LLM) and the vision encoder are frozen, while the MLP remains trainable. The learning rate is set to $1e^{-3}$ with a cosine decay schedule, using AdamW as the optimizer. Input size is $336\times 336$, and the total batch size is $32\times 8\times 2$. Here, 32 refers to the number of GPUs used, 8 represents the micro-batch size per GPU, and 2 is the number of gradient accumulation steps. Notably, we use a soft packing technique, which allows each sample to actually contain multiple sequences—estimated at over 2 per sample.
  • Stage II (Deep alignment). The LLM remains frozen, while the vision encoder and MLP are unfrozen and trainable, allowing for deeper feature alignment. The learning rate is lowered to $1e^{-4}$, the total batch size is $32\times 4\times 4$, and other settings remain consistent with Stage I.
  • Stage III (Instruction tuning). In this stage, all components of the model are trainable, allowing for comprehensive fine-tuning. The learning rate is further reduced to $1e^{-5}$, the total batch size is $32\times 4\times 4$, while other parameters, including optimizer, remain unchanged across stages.
Each stage utilizes DeepSpeed for mixed-precision training (bf16) on 32 A100 (80GB) GPUs. To accelerate the training rate of the model, we utilize the soft packing method in xtuner***https://github.com/InternLM/xtuner, which packs multiple sequences into each sample to optimize memory usage and efficiency..
Figure 7: Examples of our model in various medical tasks, including image description, disease diagnosis, diagnostic report generation, free question answering, etc. The results demonstrate the model’s versatility across different imaging modalities and tasks, with support for both Chinese and English.
11 Results
Fig. 7 illustrates several examples of our model on various tasks including image description, disease diagnosis, diagnostic report generation, free question answering, etc. These examples cover a diverse range of medical imaging modalities and tasks, highlighting the model’s versatility and general capabilities within medical domain. Furthermore, GMAI-VL supports both Chinese and English, thereby facilitating bilingual functionality and expanding its applicability in diverse clinical settings.

### v2

@@
 Abstract
-Despite significant advancements in general artificial intelligence, such as GPT-4, their effectiveness in the medical domain (general medical AI, GMAI) remains constrained due to the absence of specialized medical knowledge.
-To address this challenge, we present GMAI-VL-5.5M, a comprehensive multimodal medical dataset created by converting hundreds of specialized medical datasets into meticulously constructed image-text pairs.
-This dataset features comprehensive task coverage, diverse modalities, and high-quality image-text data.
-Building upon this multimodal dataset, we propose GMAI-VL, a general medical vision-language model with a progressively three-stage training strategy.
-This approach significantly enhances the model’s ability by integrating visual and textual information, thereby improving its ability to process multimodal data and support accurate diagnosis and clinical decision-making.
-Experimental evaluations demonstrate that GMAI-VL achieves state-of-the-art results across a wide range of multimodal medical tasks, such as visual question answering and medical image diagnosis.
-Our contributions include the development of the GMAI-VL-5.5M dataset, the introduction of the GMAI-VL model, and the establishment of new benchmarks in multiple medical domains.
-Code and dataset will be released at https://github.com/uni-medical/GMAI-VL.
+Despite significant advancements in general AI, its effectiveness in the medical domain is limited by the lack of specialized medical knowledge.
+To address this, we formulate GMAI-VL-5.5M, a multimodal medical dataset created by converting hundreds of specialized medical datasets with various annotations into high-quality image-text pairs.
+This dataset offers comprehensive task coverage, diverse modalities, and rich image-text data.
+Building upon this dataset, we develop GMAI-VL, a general medical vision-language model, with a three-stage training strategy that enhances the integration of visual and textual information.
+This approach significantly improves the model’s ability to process multimodal data, supporting accurate diagnoses and clinical decision-making.
+Experiments show that GMAI-VL achieves state-of-the-art performance across various multimodal medical tasks, including visual question answering and medical image diagnosis.
 1 Introduction
-Large-scale Vision-Language Models (LVLMs) have rapidly evolved in recent years, effectively integrating visual perception with language understanding by leveraging large-scale multimodal data, which enables them to capture complex visual and textual patterns and drive significant advancements in image recognition, natural language processing, and multimodal tasks.
-With the advancement of multimodal integration technologies, the demand for high-precision processing of diverse data types in the medical field has become increasingly critical.
-The ability to effectively integrate and analyze various data modalities, such as medical images, clinical text, and structured clinical records, is pivotal for achieving accurate and comprehensive diagnostic and treatment decisions.
-However, existing LVLMs, such as GPT-4 [1], are limited in medical applications due to their lack of domain-specific knowledge, highlighting the need for specialized solutions that effectively integrate medical expertise.
-Addressing this challenge requires constructing a comprehensive medical vision-language dataset and developing domain-specific models.
-For the medical dataset, it should provide high-quality medical knowledge, including the following three aspects:
-Comprehensive medical tasks.
-To enhance the model’s applicability across various medical scenarios, the dataset should cover a wide range of medical contexts, such as disease types, symptoms, and treatments.
-Comprehensive task coverage can improve the model’s generalization ability and increase its reliability in real-world applications.
-However, existing models often focus on specific domains [28, 41, 72, 65, 35], limiting their broader applicability.
-Expanding the dataset’s scope will further enhance the model’s utility in clinical practice.
-Rich multimodal representation.
-A well-rounded medical multimodal dataset should encompass various modalities, including different medical imaging types (such as CT, MRI, and X-rays) and diverse forms of textual data (such as medical records and imaging reports).
-This would allow models to better integrate multi-source information and improve their analytical capabilities.
-However, existing methods tend to focus on a single type of medical imagery [33, 70, 50], limiting the model’s adaptability to diverse clinical scenarios.
-A more diverse multimodal dataset would provide a foundation for developing more comprehensive models, better suited to the complexity of real-world medical environments.
+Recent advancements in Large-scale Vision-Language Models (LVLMs) have driven progress in image recognition, natural language processing, and multimodal tasks, leveraging the power of multimodal datasets.
+In the medical field (general medical AI, GMAI), as these technologies mature, the need for accurate processing of diverse data—such as medical images, clinical text, and structured records—has become critical for reliable diagnostic and treatment decisions.
+However, existing LVLMs, like GPT-4o [1], face limitations in medical applications due to their lack of domain-specific knowledge.
+This highlights the need for specialized solutions that incorporate medical expertise.
+To address this, we develop a comprehensive medical vision-language dataset and corresponding domain-specific models.
+As shown in Fig.
+1 (a), Our dataset provides high-quality medical knowledge across three aspects: (i) Comprehensive medical tasks: To improve the model’s applicability in various medical scenarios, our dataset covers a wide range of disease types, symptoms, treatments, and medical workflows.
+(ii) Rich multimodal representation: Our dataset includes a variety of modalities, such as different types of medical images (e.g., CT, MRI, X-rays) and textual data (e.g., medical records, imaging reports).
+(iii) High-quality image-text data: The performance of medical LVLMs heavily relies on high-quality image-text pairs.
+We use the well-curated collection of medical images paired with precise textual descriptions to build our dataset.
 Figure 1: Overview of GMAI-VL and GMAI-VL-5.5M.
-(a) illustrates the sources, departments, modalities, task types, and instruction formats of the GMAI-VL-5.5M dataset.
-(b) Architecture of GMAI-VL, integrating a Vision Encoder, Projector, and Large Language Model.
-(c) Three-stage training process of GMAI-VL, including shallow alignment, deep alignment, and instruction tuning with corresponding data sizes and training components.
-The flame symbol Refer to caption denotes the training part, while the snowflake symbol Refer to caption indicates frozen part.
-High-quality image-text data.
-High-quality training data is essential for optimizing model performance, especially in medical applications.
-Ideally, image-text data for medical contexts should consist of a large, well-curated collection of medical images paired with precise textual descriptions.
-This data would enhance the model’s understanding of key medical concepts, such as diagnosis, treatment, and clinical workflows, ultimately contributing to improved clinical outcomes.
-While progress has been made with datasets sourced from platforms like PubMed [39, 54, 70], these datasets suffer from inconsistent data quality, imprecise alignment between images and text, and a lack of standardization, which limits their effectiveness.
-Based on above observations, we propose an annotation-guided data generation methodology for developing a comprehensive multimodal medical datasets.
-The methodology begins by collecting large-scale, open-source medical imaging datasets and extracting key annotations, such as modality, task type, labels, and bounding boxes.
-A vision-language model (e.g., GPT-4o) is then used to transform these datasets, covering tasks like lesion detection, segmentation, and disease diagnosis, into high-quality image-text pairs for training LVLMs.
-To ensure data quality, extracted image information is incorporated into the prompt design, improving model performance across various clinical tasks.
-This results in a comprehensive multimodal dataset with 5.5M samples, named GMAI-VL-5.5M, which supports the development of general medical LVLMs.
-Fig.
-1(a) illustrates the sources, departments, modalities, task types, and instruction formats of the constructed dataset.
-With the constructed GMAI-VL-5.5M dataset, we develop a general medical vision-language model, GMAI-VL.
-To enhance its integration of visual and linguistic features and its instruction-following abilities, a three-stage training strategy is proposed in this paper.
-Specifically, we sequentially implement shallow and deep alignments in the first two stages, gradually building associations between visual (medical images) and language (medical texts) elements from basic features to high-level semantics.
-Next, we fine-tune the model with cross-modal instructions, improving its understanding of visual-language interactions and instruction-following in complex tasks.
-With this strategy, GMAI-VL shows strong performance in medical tasks like visual question answering and medical image diagnosis, providing a solid foundation for advancing multimodal models in the medical field.
+(a) Sources, departments, modalities, task types, and instruction formats of GMAI-VL-5.5M.
+(b) Architecture of GMAI-VL, with a Vision Encoder, Projector, and Large Language Model.
+(c) Three-stage training process, including shallow alignment, deep alignment, and instruction tuning, with corresponding data sizes and components.
+Refer to caption indicates the training part while Refer to caption indicates the frozen part.
+To achieve this, we develop an annotation-guided data generation process to create this large-scale multimodal medical dataset.
+First, we collect multiple open-source medical imaging datasets and extract their key annotations (e.g., modality, task type, labels, bounding boxes).
+Next, we convert these well-labeled annotations into text descriptions suitable for training vision-language models.
+This ensures that the medical image-text pairs are accurately constructed, with annotations verified by medical experts in the collected medical imaging datasets.
+The process results in the GMAI-VL-5.5M dataset, consisting of 5.5 million samples, which supports the development of general medical LVLMs.
+Building upon this dataset, we develop a general medical vision-language model, GMAI-VL.
+To optimize its ability to integrate visual and linguistic features and follow complex instructions, we present a three-stage training strategy.
+The first two stages involve shallow and deep alignments, gradually establishing connections between medical images and texts, from basic features to high-level semantics.
+In the final stage, the model is fine-tuned with cross-modal instructions, improving its understanding of visual-language interactions and its ability to follow intricate tasks, as shown in Fig.
+1 (b)&(c).
+We conduct experiments on various multimodal medical benchmarks [39, 30, 44, 77, 16, 32, 75], and the results demonstrate that GMAI-VL significantly outperforms existing models.
+Specifically, GMAI-VL sets new benchmarks with an average score of 88.48% on OmniMedVQA, 62.43% on the GMAI-MMBench test set, and 51.3% on the Health & Medicine track of MMMU.
 The contributions of this work are as follows:
-  • We build an annotation-guided data generation methodology to construct GMAI-VL-5.5M, a comprehensive vision-language dataset.
-This dataset offers extensive coverage of medical tasks, diverse multimodal representations, and high-quality image-text pairs, providing a robust foundation for model training.
-  • Leveraging GMAI-VL-5.5M, we design GMAI-VL, a versatile medical vision-language model.
-Our proposed three-stage training strategy enhances its ability to integrate visual and linguistic features, significantly improving its performance in instruction-following and generalization across a wide range of medical tasks.
-  • GMAI-VL outperforms existing models in multimodal question-answering tasks, including PMC-VQA and VQA-RAD.
-It sets new benchmarks on OmniMedVQA, GMAI-MMBench, and Health & Medicine track of MMMU.
-Specifically, GMAI-VL achieves an average score of 88.48% on OmniMedVQA, 62.43% on the GMAI-MMBench test set, and 51.3% on the Health & Medicine track of MMMU.
+  • We develop an annotation-guided data generation methodology to create GMAI-VL-5.5M, a comprehensive medical vision-language dataset with diverse medical tasks, rich multimodal representations, and high-quality image-text pairs.
+  • Leveraging this dataset, we design GMAI-VL, a general medical vision-language model, and propose a three-stage training strategy to enhance its integration of visual and linguistic features, improving its performance in various medical tasks.
+  • We demonstrate that GMAI-VL outperforms existing models in multimodal question-answering tasks, setting new benchmarks with high performance on OmniMedVQA, GMAI-MMBench, and the Health & Medicine track of MMMU.
 2 Related Work
-Large-scale medical vision-language datasets of high quality and multiple modalities are the basis of Large Vision-Language Models (LVLMs) in the medical domain.
-While natural language and vision datasets are easily accessible online, biomedical datasets often focus on text or images only and many of them are limited to specific tasks or modalities, thus with unsatisfactory generalization ability.
-Notable datasets like MIMIC-CXR [33] and CheXpert [12] have advanced radiology models but are restricted to single image modality (X-ray), which hinders their use as general-purpose medical LVLMs.
-To address this, researchers have begun scraping public sources like PubMed and textbooks to construct large-scale vision-language datasets.
-Examples include datasets proposed in LLaVA-Med [39], Med-Flamingo [54], and PubMedVision [14], with PubMedVision optimizing LLaVA-Med dataset for higher-quality medical data.
-In addition to scraping, open-source image datasets with annotations can also be converted into image-text pairs for model training.
-Specifically, image information like modalities and annotations are input into large language models, e.g., GPT series, to generate text paired with the corresponding image.
-Some popular examples include the datasets constructed in RadFM [70], MedDr [28], MedTrinity-25M [71], ChiMed-VL [48], BiomedGPT [79], Med-Gemini [58], and Med-PaLM [59].
-These efforts usually suffer from either limited modalities, data sources, or task coverage.
-Thus, their dataset quality needs further improvement.
-To this end, we propose to construct a comprehensive medical vision-language dataset with extensive coverage of medical tasks, diverse multimodal representations, and high-quality image-text pairs, forming a robust foundation for model training.
-Medical vision-language models
-are usually based on general-purpose Large Vision-Language Models (LVLMs).
-Most of them adapt LVLMs to specific medical applications using specialized medical datasets.
-For instance, Med-Flamingo [54] enhances OpenFlamingo-9B using 0.8 million interleaved and 1.6 million paired medical image-text data, highlighting the critical need for multimodal data in medical image analysis and automated report generation tasks.
-RadFM [70] improves PMC-LLaMA [69] by leveraging 16 million radiology images with text descriptions from diverse sources.
-Similarly, Med-PaLM [68] adapts PaLM-E [25] to the medical domain with approximately one million medical data samples, achieving state-of-the-art performance in diagnostic support and medical knowledge Q&A.
-LLaVA-Med [39] utilizes a large-scale biomedical figure-caption dataset extracted from PubMed Central to enhance LLaVA [66, 67] to better understand biomedical images and facilitate open-ended conversational interactions.
-Med-Gemini [58] leverages long-format question-answering datasets to improve the multimodal and long-contextual capabilities of the baseline Gemini model, enabling superior performance in complex medical Q&A and multimodal reasoning tasks.
-Additionally, HuatuoGPT-Vision [14] and MedDr [28] build medical multimodal datasets to adapt general-purpose LVLMs like LLaVA and InternVL to various medical modalities, including radiology, pathology, dermatology, and endoscopy.
-Previous studies usually focus on constructing medical datasets to adapt general-purpose LVLMs but pay less attention to the adaptation strategies.
-However, naive training/adaptation strategies may not successfully adapt general-purpose LVLMs to the medical data, due to the large gap between the natural image-text pairs and the medical ones.
-Moreover, these strategies can hardly align the broad imaging modalities and various types of medical text (e.g., prescriptions, radiology reports, and electronic health records) to obtain generalizable features, thus limiting the models’ performance.
-Our work thus proposes a novel three-stage training strategy to better integrate the visual and language features to enhance generalization ability.
-3 GMAI-VL-5.5M: A Comprehensive Multimodal Dataset
-With rapid advancements in medical vision-language models (VLMs), constructing high-quality datasets is crucial for developing general-purpose models.
-Unlike previous approaches that primarily rely on published literature, our method utilizes specialized medical datasets to create a more robust, high-quality resource.
-We build GMAI-VL-5.5M, a comprehensive medical vision-language dataset that aggregates data from diverse open-source and proprietary sources.
-Covering 13 medical imaging modalities and 18 specialties, it addresses a wide range of medical imaging tasks.
-This dataset enhances the model’s ability to process complex medical information, advancing precision medicine and intelligent diagnostics.
-Table 1: Comparison of various medical multimodal datasets, including details on the dataset size, modality type, language, data traceability, and sources of information.
+Large-scale medical vision-language datasets.
+These datasets are crucial for training Large Vision-Language Models (LVLMs) in the medical domain.
+While general datasets are readily available, biomedical datasets often focus on text or images separately, limiting their generalization.
+Datasets like MIMIC-CXR [35] and CheXpert [12] advance radiology models but are restricted to a single image modality (X-ray), limiting their use as general-purpose medical LVLMs.
+To address this, researchers have compiled large-scale vision-language datasets by scraping public resources like PubMed and medical textbooks.
+Datasets like LLaVA-Med [40], Med-Flamingo [53], and PubMedVision [14] improve upon LLaVA-Med by enhancing the quality of medical data.
+Additionally, open-source image datasets with annotations have been converted into image-text pairs for training.
+Notable examples include RadFM [68], MedDr [29], BiomedGPT [76], Med-Gemini [57], and Med-PaLM [58].
+MedTrinity-25M [69] generates image-text pairs for supervised fine-tuning.
+However, these datasets often face limitations in modalities, data sources, or task coverage, necessitating improvements.
+We construct a comprehensive medical vision-language dataset with broad task coverage, diverse modalities, and high-quality image-text pairs to provide a solid foundation for model training.
+Table 1: Comparison of multimodal medical datasets, including size, modality, language, traceability, and sources.
 Datasets Data Size Modality Language Traceability Data Source
@@
 GMAI-VL-5.5M (ours) 5.5M Multimodal EN&CN 219 specialized medical imaging datasets
+Medical vision-language models.
+Medical vision-language models often adapt general-purpose Large Vision-Language Models (LVLMs) for specific medical tasks using specialized datasets.
+For example, Med-Flamingo [53] improves OpenFlamingo-9B with 0.8 million interleaved and 1.6 million paired medical image-text data for medical image analysis and report generation.
+RadFM [68] enhances PMC-LLaMA [67] with 16 million radiology images and text descriptions.
+Med-PaLM [66] adapts PaLM-E [25] to medical data, achieving state-of-the-art results in diagnostic support and Q&A.
+LLaVA-Med [40] uses a biomedical figure-caption dataset from PubMed Central to improve LLaVA [64, 65] for biomedical image understanding and open-ended conversations.
+Med-Gemini [57] leverages long-format question-answering datasets for better multimodal and contextual capabilities, enhancing complex medical Q&A and reasoning tasks.
+Additionally, HuatuoGPT-Vision [14] and MedDr [29] adapt general-purpose LVLMs like LLaVA and InternVL to various medical modalities, including radiology, pathology, dermatology, and endoscopy.
+While these studies focus on constructing medical datasets, they often overlook adaptation strategies.
+Naive training methods may fail to bridge the gap between natural and medical image-text pairs, or align diverse medical modalities and texts (e.g., prescriptions, radiology reports, EHRs), limiting generalizability.
+Our work introduces a novel three-stage training strategy to better integrate visual and language features, improving generalization.
+Table 2: The examples of annotation-guided prompts for paired data format generation.
+Instruction-following data prompt
+[Uncaptioned image] As a medical expert, you will receive a <fundus>image belonging to <Ophthalmology>, which is labeled as <moderate nonproliferative diabetic retinopathy>.
+Please generate question-answer pairs based on the following requirements: 1.
+The question-answer pairs must be strictly based on the content of the image and directly related to the label information, should not state known medical facts or general definitions directly.
+2.
+Focus on reasoning around the label and guide observation of the visible features of the disease, such as affected areas, symptoms, and changes in the retina.
+3.
+Each question and answer must be precise, clear, and derived from observable characteristics in the image, with no vague or uncertain information.
+4.
+The output format must follow the standard JSON format with the structure as follows: { ’question’: <question content >, ’answer’: <answer content >}
+Question Can we see any hard exudates or fluid leakage in the image?
+Answer The image does show some bright, yellowish deposits, which are consistent with hard exudates, often seen in moderate nonproliferative diabetic retinopathy due to fluid leakage.
+Question What is the appearance of the optic disc in this retinal image?
+Answer The optic disc appears normal without any visible signs of swelling or abnormal changes, suggesting that the optic nerve is not yet affected by the disease.
+Caption data prompt
+[Uncaptioned image] As a medical expert, you are given a <X-ray> image belonging to <Pulmonary Medicine>, where the <bbox [179,164,500,434]> is labeled as <Pneumothorax>.
+Please generate a very detailed description related to bbox in the given image.
+When generating the description, you must strictly adhere to the following guidelines: 1.
+The generated description should be inferable from the image; 2.
+The generated description should include image features related to the given label; 3.
+The generated description should start with ’This region’; 4.
+The description may imply known information but should not directly state known information; 5.
+The generated description should include major entities within and around the region, background elements, texture features, and any other notable features.
+Output: This region shows … the right hemithorax on a frontal chest X-ray, … an area suggestive of abnormal air accumulation.
+The bounded area appears less dense compared to surrounding lung tissue, … The edges of the lung are visible, separated from the chest wall by a faint radiolucent (darker) space, suggestive of pneumothorax.
+The lung markings are notably absent within this area, … The texture within the defined area has a homogeneous, darkened appearance consistent with air rather than tissue density, contrasting with the granular pattern in the aerated lung parenchyma outside this region.
+3 GMAI-VL-5.5M: A Comprehensive Multimodal Medical Dataset
+With rapid advancements in medical vision-language models, constructing high-quality datasets is essential for developing general-purpose models.
+Unlike previous approaches that mainly rely on published literature, our method leverages specialized medical datasets with various annotations to create a more robust, high-quality resource.
+We present GMAI-VL-5.5M, a comprehensive medical vision-language dataset that aggregates data from diverse open-source and proprietary sources.
+Covering 13 medical imaging modalities and 18 specialties, it supports a wide range of medical imaging tasks.
+This dataset enhances the model’s ability to process complex medical information, advancing precision medicine and intelligent diagnostics.
 3.1 Data Curation
@@
 They also encompass multiple clinical specialties, including pathology, dermatology, ophthalmology, otolaryngology, and oncology, further enhancing their diversity.
-Data preprocess.
-After data collection, we apply a preprocessing workflow to extract 2D medical images from the videos and 3D medical volumes.
-For data cleaning, we first extract key annotation information from classification data, including modality, department, and labels for each image, discarding instances with missing or unclear annotations.
-For segmentation data, we follow the SA-Med2D-20M [75] preprocessing method, filtering out low-quality images and labels, and converting them into detection datasets.
-The converted segmentation data is then processed together with the original detection dataset, with key annotation information extracted and linked to each image.
-Finally, the preprocessed data are then standardized and organized into a structured format: <image, modality, label, department, bbox [optional]>.
-Data generation.
-Subsequently, the data are categorized into two primary types: classification datasets and detection/segmentation datasets.
-Each category is further refined using specific prompts tailored for large model training.
-For data generation, large vision-language models (GPT-4o) are employed to produce detailed image descriptions and corresponding instruction-following data through annotation-guided data generation methodology.
-For classification datasets, detailed descriptions of the entire image are generated, while for detection datasets, the focus is on specific regions enclosed by bounding boxes, providing comprehensive functional analyses of these areas.
-Notably, the segmentation dataset was transformed into a detection dataset using external bounding boxes, and data generation followed detection dataset protocols.
-Furthermore, to improve the model’s multilingual capability, we translated a portion of English image-text data into Chinese.
-Incorporating multilingual data helps to enhance the generalization capabilities of domain-specific multimodal models.
-The resulted data is utilized for medical Visual Question Answering (VQA) tasks, forming the comprehensive VQA dataset, named GMAI-VL-5.5M.
-The detailed pipeline for generating prompt-driven data is illustrated in Fig.
-2.
-Figure 2: The prompt-driven data generation pipeline comparing without-annotation-guided and annotation-guided methods.
-The annotation-guided approach integrates specific annotation information (e.g., <image, modality, label, department, bbox [optional]>) to generate high-quality, accurate descriptions, while the without-annotation-guided approach often results in lower-quality outputs.
-Figure with complete prompt and response is provided in Supp.
-Mat..
-As depicted in Fig.
-1(a), it contains six kinds of instrcuction-following formats, including image-level captions, region-level captions, free instructions, dialogue, visual perception and text-only tasks.
-The specific composition of GMAI-VL-5.5M can be found in Supp.
-Mat.
-(Table 6).
-These formats enable VLMs to better understand and process complex visual and textual information in medical contexts.
-The GMAI-VL-5.5M dataset significantly enhances the model’s cross-modal reasoning ability, enabling it to handle complex multimodal inputs in real clinical scenarios.
-The richness of the instruction formats allow the model to progress from basic question answering to advanced medical image analysis, ultimately providing strong support for clinical diagnosis and decision-making.
+Data processing.
+After data collection, we follow a workflow to pair medical data (including both 2D and 3D data) with corresponding text annotations.
+To ensure high-quality annotations, we first extract key information from the annotations provided by medical experts.
+For classification data, we extract the modality, department, and labels for each image, discarding instances with missing or unclear annotations.
+For segmentation data, we follow the SA-Med2D-20M [72] approach, filtering out low-quality images and labels, and converting them into detection annotation format.
+Finally, the preprocessed data is standardized and organized into a structured format: <image, modality, label, department, bbox [optional]>, where “bbox” refers to the bounding box locations for detection annotations.
+Paired data format generation.
+To generate paired visual medical data with text descriptions, we use large vision-language models (i.e., GPT-4o) to produce detailed descriptions with instructions via an annotation-guided methodology.
+In details, for classification datasets, comprehensive descriptions of the entire image are created, while for detection datasets, the focus is on specific regions enclosed by bounding boxes, with detailed functional analyses of these areas.
+Furthermore, based on the given information, instruction-following question-answer pairs related to the medical images are generated.
+These pairs involve specific instructions tailored to the medical context, such as identifying critical anatomical or pathological features within the medical images (e.g., tumors, lesions, or organs) or providing in-depth interpretations of regions of interest.
+These instructions guide the model to produce relevant, precise responses, thereby enhancing its applicability to specialized medical imaging tasks.
+As mentioned in the previous subsection, the segmentation dataset is transformed into a detection format using external bounding boxes, and data generation follows the detection dataset protocols.
+The detailed example prompts are shown in Table 2.
+The generated data is then used for medical Visual Question Answering (VQA) tasks, forming the comprehensive VQA dataset, GMAI-VL-5.5M.
+To enhance the model’s multilingual capability, we translate approximately 30% of the English image-text data into Chinese.
+This multilingual data helps further improve the generalization ability of domain-specific multimodal models.
 3.2 Data Property
 Data statistics.
-These datasets encompass diverse medical imaging tasks and modalities, forming a solid foundation for developing and evaluating medical LVLMs.
-Fig.
-4 (in Supp.
-Mat.) illustrates the distribution of modalities, tasks, clinical departments, and specific medical challenges represented within the collected datasets.
-This visualization highlights the extensive diversity and coverage of our data collection efforts.
-After careful standardization and integration, these datasets form the core of our comprehensive medical image-text dataset, GMAI-VL-5.5M, which serves as a crucial resource for advancing precision medicine and intelligent diagnostic systems.
-Accuracy and reliability of generated data.
-The accuracy and reliability of the generated data are primarily addressed from two key aspects.
-First, the reliability of the data source is ensured by utilizing high-quality datasets from reputable sources.
-The data used for generation primarily comes from professional medical challenges (such as Kaggle and GrandChallenge) and publicly available, peer-reviewed datasets.
-This guarantees the integrity and accuracy of the data at its source.
-Second, the data generation process is rigorously controlled to ensure accuracy and reliability.
-Although GPT was used for data generation, prompts were meticulously crafted to include essential annotations, minimizing errors and hallucinations.
-As illustrated in Fig.
-2, the proposed annotation-guided approach incorporates specific annotations (e.g., <image, modality, label, department, bbox [optional]>) to produce precise, high-quality descriptions.
-Compared to the without-annotation-guided approach, which often results in less detailed outputs, the annotation-guided method delivers more professional, pathology-specific descriptions, significantly enhancing the reliability of the generated data.
+Our dataset encompasses a broad spectrum of medical imaging tasks and modalities, forming a robust foundation for the development and evaluation of medical LVLMs.
+Table 3 summarizes the distribution of key modalities, tasks, clinical departments, and clinical tasks within GMAI-VL-5.5M, illustrating its diversity and extensive coverage.
+For a more detailed analysis, additional visual insights into the dataset composition can be found in the supplementary material.
+Data quality.
+The quality of the generated paired data and annotations is ensured through two main approaches.
+First, we use high-quality datasets from trusted sources, including professional challenges like Kaggle and GrandChallenge, as well as peer-reviewed datasets, ensuring data accuracy and reliability.
+Second, we carefully control the data generation process.
+While GPT is used for generation, the prompts are designed with essential annotations (e.g., <image, modality, label, department, bbox [optional]>) to minimize errors.
+This annotation-guided approach produces more detailed and professional descriptions.
+The data quality is detailed in the supplementary material.
+Table 3: Distribution of GMAI-VL-5.5M across key dimensions.
+Dimension Value Percentage
+Task Type 2D Classification 50.4%
+3D Segmentation 30.3%
+2D Segmentation 12.7%
+2D Detection 6.6%
+Modality CT 26.8%
+MR 24.7%
+Endoscopy 12.6%
+Pathology 11.2%
+X-Ray 6.7%
+Fundus 5.3%
+Ultrasound 3.0%
+Department Orthopedic Surgery 12.9%
+General Surgery 10.3%
+Gastroenterology 9.7%
+Hematology 9.2%
+Pulmonary Medicine 9.0%
+Sports Medicine 8.2%
+Clinical Task Disease Diagnosis 40.4%
+Organ Recognition 16.0%
+Bone Recognition 8.5%
+Severity Grading 6.1%
+Surgeon Action Recognition 6.0%
 Compared with other medical multimodal datasets.
-The GMAI-VL-5.5M dataset, as highlighted in Table 1, stands out due to its unmatched scale, encompassing over 5.5 million samples from more than 219 specialized medical imaging datasets.
-Unlike other datasets listed, GMAI-VL-5.5M supports a wider variety of modalities and languages, making it a truly global resource that caters to diverse clinical needs.
-Additionally, GMAI-VL-5.5M emphasizes traceability of its data, ensuring a high standard of clinical relevance and reliability.
-This comprehensive and diverse dataset is critical for pushing the boundaries of medical multimodal research, enabling more effective training of LVLMs that can generalize across multiple medical tasks and scenarios, thereby driving innovations in precision medicine and intelligent diagnostics.
+The GMAI-VL-5.5M dataset, as shown in Table 1, distinguishes itself with its unmatched scale, featuring over 5.5 million samples from more than 219 specialized medical imaging datasets.
+Unlike other datasets, GMAI-VL-5.5M supports a broader range of modalities and languages, making it a truly global resource that addresses diverse clinical needs.
+Furthermore, GMAI-VL-5.5M ensures data traceability, maintaining high clinical relevance and reliability.
+This extensive and varied dataset is crucial for advancing medical multimodal research, enabling more effective training of LVLMs that can generalize across numerous medical tasks and scenarios, ultimately driving innovations in precision medicine and intelligent diagnostics.
 4 GMAI-VL: A General Medical Vision-Language Model
@@
 1(b).
-These components are designed to work together seamlessly, enabling the model to deliver exceptional performance in medical applications.
-We adopt InternLM2.5-7B [63] as our language processing module, which offers outstanding reasoning capabilities.
-With a context window up to one million tokens, it can handle complex medical tasks and generate coherent, accurate responses.
-Its support for advanced instruction-following makes it particularly effective in addressing intricate medical queries, thereby enhancing the model’s ability to understand and respond to a wide range of instructions.
-For vision processing, GMAI-VL employs a CLIP-based vision encoder [55], which transforms visual inputs into high-dimensional feature representations.
-CLIP’s strong performance in aligning image and text representations ensures that medical image features are accurately extracted and effectively integrated with linguistic information, significantly enhancing the model’s ability to handle multimodal medical data.
-The MLP, as a projector, serves as a bridge between the vision encoder and the LLM, optimizing high-dimensional outputs and further enhancing feature representation.
-The seamless integration of these components enables GMAI-VL to excel in processing and understanding complex medical multimodal data.
-4.2 Optimization Strategy
-As illustrated in Fig.
-1(c), the training process of the GMAI-VL model is divided into three stages: shadow alignment, deep alignment, instruction tuning, respectively.
-The detailed hyper-parameter settings can be found in Supp.
-Mat.
-(Table 8).
-To enhance the training of GMAI-VL, we supplement our GMAI-VL-5.5M dataset with additional medical datasets, ultimately forming an 11.7M dataset for training.
-This supplemented data increases the diversity of training data, exposing the model to a wider range of medical scenarios and visual-language patterns, enhancing its generalization to complex clinical tasks and ensuring robustness in real-world applications.
-Fig.
-5 (in Supp.
-Mat.) provides a complete distribution of the utilized datasets during the training stage.
-The detailed data proportions for each training stage are detailed in Table 7 (in Supp.
-Mat.).
+We use InternLM2.5-7B [62] as our language processing module, which provides exceptional reasoning capabilities.
+With a context window of up to one million tokens, it can manage complex medical tasks and generate coherent, accurate responses.
+For vision processing, we adopt a CLIP-based vision encoder [54], which converts visual inputs into high-dimensional feature representations.
+The MLP projector bridges the vision encoder and language processing module, optimizing high-dimensional outputs and improving feature representation.
+This framework enables GMAI-VL to effectively process multimodal medical data.
+4.2 Training Strategy
+As shown in Fig.1(c), the training process of the GMAI-VL model is divided into three stages: shallow alignment, deep alignment, and instruction tuning.
+To enhance the training of GMAI-VL, we supplement the GMAI-VL-5.5M dataset with additional medical datasets, resulting in a total training dataset of 11.7M samples.
 Stage I: Shadow alignment.
-In the shallow alignment phase, we utilize a large-scale medical image-text dataset comprising approximately 11.7 million image-text pairs, sourced from a combination of publicly accessible datasets and proprietary in-house data.
-To achieve shallow alignment, we freeze both the large language model and the vision encoder, optimizing only the projector.
-With this optimization stage, the model establishes an initial alignment between medical images and their corresponding textual descriptions.
-All input images are resized to $336\times 336$ pixels, and the training objective is to minimize the cross-entropy loss of the text tokens.
+In this phase, we use a large-scale medical image-text dataset of approximately 11.7 million image-text pairs.
+During this stage, we freeze both the large language model and the vision encoder, optimizing only the projector.
+This optimization establishes an initial alignment between medical images and their corresponding textual descriptions.
+During training, (i) the objective is to minimize the cross-entropy loss of the text tokens; (ii) all images are resized to $336\times 336$ pixels; (iii) the learning rate is set to $1e^{-3}$ with a cosine decay schedule, and AdamW is used as the optimizer; (iv) the total batch size is $32\times 8\times 2$, where 32 refers to the number of GPUs used, eight represents the micro-batch size per GPU, and two is the number of gradient accumulation steps; and (v) a soft packing technique is used to allow each sample to contain multiple sequences, averaging over two sequences per sample.
 Stage II: Deep alignment.
-Most vision encoders in multimodal models are pre-trained on natural images, so it is essential to address the domain gap between medical and natural images during the deep alignment stage.
-To achieve this, we fine-tune both the vision-language projector and the vision encoder, enhancing the alignment between the visual features of medical images and the language model’s feature space.
+Most vision encoders in multimodal models are pre-trained on natural images, so addressing the domain gap between medical and natural images is critical during the deep alignment stage.
+To bridge this gap, we fine-tune both the vision-language projector and the vision encoder, enhancing the alignment between the visual features of medical images and the language model’s feature space.
+The learning rate is set to $1e^{-4}$, the total batch size is $32\times 4\times 4$, and other settings remain consistent with Stage I.
 Stage III: Instruction tuning.
-At this stage, we fine-tune our GMAI-VL model (including the vision encoder, the language model, and the projector components) by instruction tuning to enhance its instruction-following and dialogue capabilities.
-The multimodal instruction tuning data is primarily derived from the training data in previous stages, with a filtering process to select high-quality and more suitable data for fine-tuning.
-Specifically, we manually reviewed a subset of each dataset and labeled it as either “high quality” or “low quality.” High-quality datasets were used for tuning, while most low-quality ones were excluded, except for a small portion retained for diversity.
-This approach allowed efficient filtering without reviewing every sample.
-Additionally, we incorporate medical text dialogue data to ensure the model’s versatility in handling various dialogue scenarios.
-Thus, our instruction tuning data comprises approximately ten million samples.
-5 Experimental Results
-To evaluate our model, we employed several established multimodal medical benchmarks, each targeting specific aspects of medical image understanding and question answering.
-Below is a brief overview of the benchmarks used in our experiments:
-  • Traditional medical VQA benchmarks: Traditional multimodal medical question-answering benchmarks, such as VQA-RAD [38], SLAKE [44], and PMCVQA [80], span various imaging modalities tasks, primarily assessing the model’s ability to extract information from medical images and answer clinical questions.
-They evaluate the model’s performance in understanding medical imaging and integrating multimodal information.
-  • OmniMedVQA: OmniMedVQA [31] provides a rich dataset of paired medical images and text, designed to evaluate the model’s ability to recognize and understand fundamental medical imaging concepts, focusing on cross-modal reasoning and information integration.
-  • GMAI-MMBench: GMAI-MMBench [16] focuses on assessing the model’s ability to identify fine-grained objects in complex clinical scenarios, challenging its capacity to handle long-context tasks and accurately recognize and reason over detailed medical features.
-  • MMMU Health & Medicine track: The Health & Medicine track of the MMMU [78] benchmark spans a wide range of medical fields, derived from university exams, quizzes, and textbooks.
-It evaluates the model’s reasoning ability in complex medical scenarios and the specialized knowledge in health and medicine.
-5.1 Medical VQA Benchmarks
-Table 2: Results on Traditional Medical VQA Benchmarks. The highest performance in each column is highlighted in red, and the second-highest performance is highlighted in blue.
+In this stage, we fine-tune the entire GMAI-VL model—vision encoder, language model, and projector—through instruction tuning to enhance its instruction-following and dialogue capabilities.
+The multimodal instruction data is primarily derived from the previous stages.
+Note that this stage discards the low-quality samples that with too short descriptions or extremely lower confidence in the previous stages.
+Additionally, medical text dialogue data is incorporated to improve the model’s handling of various dialogue scenarios.
+This results in ten million samples for instruction tuning.
+During taining, the learning rate is set to $1e^{-5}$, the total batch size is $32\times 4\times 4$, while other parameters, including optimizer, remain unchanged across stages.
+Table 4: Results on medical VQA benchmarks. The highest performance in each column is highlighted in red, and the second-highest performance is highlighted in blue.
 Model VQA-RAD SLAKE PMC-VQA Avg.
@@
 GMAI-VL (ours) 66.3 72.9 54.3 64.5
-The performance of various VLMs on popular medical VQA benchmark datasets is summarized in Table 2, including VQA-RAD[38], SLAKE[44], and PMC-VQA[80].
-Our model, GMAI-VL, demonstrates a strong performance, achieving the highest score on the VQA-RAD[38] dataset with 66.3%, outperforming other models such as HuatuoGPT-Vision-7B.
-This result highlights GMAI-VL’s superior capability in handling radiological image question-answering tasks.
-For the PMC-VQA[80] dataset, GMAI-VL achieves 54.3%, and 72.9% on SLAKE[44], demonstrating its capability in handling medical VQA tasks across diverse modalities.
-GMAI-VL demonstrates competitive performance across multiple benchmarks, showcasing its versatility in medical image understanding and question-answering.
-5.2 OmniMedVQA
-Table 3: Comparison of performance between representative LVLMs and GMAI-VL on OmniMedVQA across five different question type.
-The best performance in each column is highlighted in red, and the second-best performance is highlighted in blue.
+5 Experimental Results
+To evaluate our model, we utilized several established multimodal medical benchmarks, including medical VQA benchmarks (PMCVQA [77], PathVQA [30], VQA-RAD [39], and SLAKE [44]) as well as benchmarks designed for large vision-language models (OmniMedVQA [32], GMAI-MMBench [16], and the MMMU Health & Medicine track [75]).
+These benchmarks target specific aspects of medical image understanding and question answering.
+Detailed information about these benchmarks can be found in the supplementary material.
+We evaluate model performance using VLMEvalKit [26] with its default settings.
+To prevent data leakage in GMAI-VL-5.5M, we ensure that the evaluation datasets, OmniMedVQA and GMAI-MMBench, are sourced exclusively from public test sets, while the training data is strictly from public training sets.
+Additionally, MD5 hashes were computed for each image to verify that there are no duplicates with the benchmark images.
+5.1 Comparisons on Medical VQA Datasets
+The performance of various VLMs on popular medical VQA benchmark datasets is summarized in Table 4.
+Our model, GMAI-VL (ours), demonstrates strong performance, achieving the highest score of 66.3% on the VQA-RAD [39] dataset, outperforming models such as HuatuoGPT-Vision-7B.
+This result underscores GMAI-VL’s superior capability in handling radiological image question-answering tasks.
+On the PMC-VQA [77] dataset, GMAI-VL achieves 54.3%, and 72.9% on SLAKE [44], highlighting its effectiveness across diverse medical VQA tasks.
+Overall, GMAI-VL demonstrates competitive performance across multiple benchmarks, showcasing its versatility in medical image understanding and question answering.
+5.2 Comparisons on OmniMedVQA
+Table 5: Comparison of LVLMs and GMAI-VL on OmniMedVQA across five question types.
+The best performance in each column is highlighted in red, and the second-best in blue.
 Abbreviations: MR = Modality Recognition, AI = Anatomy Identification, DD = Disease Diagnosis, LG = Lesion Grading, OBA = Other Biological Attributes.
@@
 GMAI-VL (ours) 98.64 92.95 88.7 87.21 82.95 88.48
-Table 3 summarizes the performance of various large vision-language models (LVLMs), including our proposed GMAI-VL, across five question types: Modality Recognition, Anatomy Identification, Disease Diagnosis, Lesion Grading, and Other Biological Attributes.
-GMAI-VL demonstrates outstanding accuracy across multiple tasks, achieving 98.64% in Modality Recognition, 92.95% in Anatomy Identification, and 88.71% in Disease Diagnosis.
-It outperforms both open-source LVLMs and medical-specific models, underscoring its capability to accurately identify anatomical structures and diagnose diseases from visual data.
-In Lesion Grading, GMAI-VL attained the highest score of 87.21%, and it also delivers a strong performance of 82.95% in Other Biological Attributes, showcasing its versatility across diverse biological contexts.
-With an average accuracy of 88.48%, the highest among all evaluated models, GMAI-VL excels not only in general medical question-answer tasks but also in complex reasoning requiring domain-specific knowledge, surpassing models like HuatuoGPT-Vision-34B and InternVL2-40B.
-These results verify our GMAI-VL is a leading model in multimodal medical image understanding, setting a new benchmark for medical VQA tasks.
-Its consistent top performance across question types highlights its potential for broader applications in medical question answering.
-5.3 GMAI-MMBench
 Table 6: Results on the val and test sets of GMAI-MMBench for clinical VQA tasks.
-The full names of the evaluated tasks can be found in Table 5 in literature [16].
-The best model in each category is highlighted in red, while the second-best model is indicated in blue.
+Full task names are in Table 5 of [16], and additional model comparisons are in Table 11 (Supplementary Materials).
+The best and second-best models are marked in red and blue, respectively.
 Model Name Overall (val) Overall (test) AR BVR B CR C DD IQG MR M NT OR-A OR-HN OR-P OR-T SG SAR SIR SWR
-Random Guess
-Random 25.70 25.94 38.20 22.73 22.92 22.72 24.06 26.66 27.13 27.00 20.00 24.75 21.37 22.93 22.33 21.18 32.43 24.23 21.39 23.71
-Open-Source LVLMs
-Flamingo v2 [4] 25.58 26.34 37.74 21.50 20.62 22.00 22.41 27.29 25.91 27.45 18.00 28.79 25.16 22.13 22.00 22.00 34.61 22.88 20.44 27.43
-VisualGLM-6B [22] 29.58 30.45 40.16 33.92 24.92 25.22 24.21 32.99 29.96 29.53 21.20 37.88 30.32 24.80 13.33 29.88 33.11 19.62 19.16 37.43
-InstructBLIP-7B  [20] 31.80 30.95 42.12 26.92 24.92 28.09 21.65 34.58 31.58 29.23 22.40 30.30 28.95 27.47 23.00 24.82 32.88 19.81 21.64 26.57
-Qwen-VL  [6] 34.80 36.05 37.05 37.24 35.85 28.98 24.81 43.60 24.70 30.12 19.20 44.44 29.68 31.87 25.00 31.18 30.26 21.54 20.10 26.86
-Yi-VL-6B  [77] 34.82 34.31 41.66 39.16 26.62 30.23 31.88 38.01 26.72 24.93 25.20 37.37 29.58 31.20 32.33 30.59 36.71 24.81 23.18 31.43
-ShareGPT4V-7B  [15] 36.71 36.70 43.96 37.59 21.54 37.57 18.80 43.26 32.39 27.30 22.80 43.43 29.47 37.33 22.00 31.76 34.98 24.42 25.06 30.00
-LLAVA-V1.5-7B [45] 38.23 37.96 45.45 34.27 30.92 41.32 21.65 44.68 34.01 27.74 23.60 43.43 28.00 42.13 29.00 35.06 33.41 22.12 23.61 29.14
-XComposer2  [24] 38.68 39.20 41.89 37.59 33.69 40.79 22.26 45.87 36.44 32.94 27.20 58.59 26.11 36.40 43.67 37.29 32.06 23.46 27.80 32.86
-LLAVA-InternLM-7b [19] 38.71 39.11 36.36 36.54 32.62 38.10 30.68 46.53 34.82 28.19 25.20 48.99 28.11 40.53 33.33 36.00 34.08 26.73 24.12 29.71
-InternVL-Chat-V1.5 [18] 38.86 39.73 43.84 44.58 34.00 33.99 31.28 45.59 33.20 38.28 32.40 42.42 31.89 42.80 27.00 36.82 34.76 23.27 24.72 32.57
-InternVL-Chat-V1.2  [17] 39.52 40.01 41.66 44.06 27.38 38.46 34.29 46.99 33.60 34.42 21.20 47.98 30.63 42.80 27.67 35.88 35.59 23.85 24.98 28.00
-LLAVA-InternLM2-7b [19] 40.07 40.45 39.82 37.94 30.62 35.24 29.77 48.97 34.01 25.96 20.80 53.03 30.95 42.67 32.00 39.88 32.43 21.73 24.38 38.00
-DeepSeek-VL-7B  [49] 41.73 43.43 38.43 47.03 42.31 37.03 26.47 51.11 33.20 31.16 26.00 44.95 36.00 58.13 36.33 47.29 34.91 18.08 25.49 39.43
-MiniCPM-V2 [73] 41.79 42.54 40.74 43.01 36.46 37.57 27.82 51.08 28.74 29.08 26.80 47.47 37.05 46.40 25.33 46.59 35.89 22.31 23.44 31.71
-Proprietary LVLMs
-Claude3-Opus [2] 32.37 32.44 1.61 39.51 34.31 31.66 12.63 39.26 28.74 30.86 22.40 37.37 25.79 41.07 29.33 33.18 31.31 21.35 23.87 4.00
-Qwen-VL-Max  [6] 41.34 42.16 32.68 44.58 31.38 40.79 10.68 50.53 32.79 44.36 29.20 51.52 41.37 58.00 30.67 41.65 26.95 25.00 24.64 39.14
-GPT-4V  [1] 42.50 44.08 29.92 48.95 44.00 37.39 12.93 52.88 32.79 44.21 32.80 63.64 39.89 54.13 37.00 50.59 27.55 23.08 25.75 37.43
-Gemini 1.0  [62] 44.38 44.93 42.12 45.10 46.46 37.57 20.45 53.29 35.22 36.94 25.20 51.01 34.74 59.60 34.00 50.00 36.64 23.65 23.87 35.43
-Gemini 1.5  [56] 47.42 48.36 43.50 56.12 51.23 47.58 2.26 55.33 38.87 48.07 30.00 76.26 51.05 75.87 46.33 62.24 20.57 27.69 30.54 40.57
-GPT-4o  [1] 53.53 53.96 38.32 61.01 57.08 49.02 46.62 61.45 46.56 56.38 34.00 75.25 53.79 69.47 48.67 65.88 33.93 22.88 29.51 39.43
 Medical Special Model
@@
 GMAI-VL (ours) 61.74 62.43 75.26 59.66 67.24 56.86 54.29 67.14 42.80 79.97 41.60 75.00 60.45 75.48 53.33 58.12 42.09 72.31 37.40 59.14
-The GMAI-MMBench benchmark is a comprehensive medical multimodal benchmark designed to evaluate models on a range of clinical visual question-answering (VQA) tasks.
-Table 4 presents the results of various LVLMs, including open-source LVLMs and commercial models, evaluated on the val and test sets across multiple clinical tasks.
-Notably, GMAI-VL excels in specific tasks such as abnormality recognition (73.78%), biological variation recognition (63.06%), and clinical disease diagnosis (66.67%).
-These results demonstrates the model’s strong ability in understanding and interpreting complex clinical images.
-In comparison to other models, GMAI-VL consistently achieves either the best or second-best performance across most tasks.
-For instance, it ranks first in 16 out of 20 categories, including key tasks such as AR (Attribute Recognition) and DD (Disease Diagnosis), where it achieved scores of 75.26% and 67.14%, respectively, suggesting GMAI-VL’s strength in understanding medical scenarios.
-Overall, GMAI-VL establishes a new benchmark in various clinical VQA tasks, demonstrating its potential as a reliable and versatile tool in medical multimodal applications.
-5.4 MMMU Health & Medicine track
-The MMMU benchmark, a widely recognized standard for evaluating multimodal models, was used to assess the performance of our proposed GMAI-VL model on the Health & Medicine track.
-The experimental results in Table 5 show the model’s performance across five key categories: Basic Medical Science (BMS), Clinical Medicine (CM), Diagnostics and Laboratory Medicine (DLM), Pharmacy (P), and Public Health (PH).
-GMAI-VL performs strongly across multiple categories, achieving top scores in DLM (43.3%), P (50.0%), and PH (53.3%), surpassing competitive models like LLaVA-v1.6 and HuatuoGPT-Vision-7B.
-These results highlight the model’s proficiency in handling complex tasks requiring diagnostic reasoning, pharmaceutical knowledge, and public health expertise.
-In BMS, GMAI-VL scores 50.0%, achieve the best performance, demonstrating the model’s the capacity of understanding medical knowledge.
-In CM, the model achieves 60.0%, remaining competitive with other leading models.
-These results underscore the model’s ability in processing both clinical and foundational medical information effectively.
-Overall, GMAI-VL achieves an average score of 51.3% across the Health & Medicine track, which is a top performance among other models, verifying its versatility in specialized medical domains.
+Table 7: Evaluation on the training strategy.
+Model MMMU_val OmniMedVQA GMAI_MMBench_val GMAI_MMBench_test
+stage1 46.00 81.38 51.49 53.69
+stage1+2 46.67 84.64 55.85 58.28
+stage1+3 45.33 86.62 59.25 60.70
+stage1+2+3 (Prop) 51.30 88.48 61.74 62.43
 Table 8: Performance on the val set for the MMMU Health & Medicine track.
-This track is divided into five categories: BMS (Basic Medical Science), CM (Clinical Medicine), DLM (Diagnostics and Laboratory Medicine), P (Pharmacy), and PH (Public Health).
-The best performance in each column is highlighted in red, and the second-best performance is highlighted in blue.
+The best performance is highlighted in red while the second-best performance is highlighted in blue.
+Note that the results are obtained from the official website.
 Model BMS CM DLM P PH MMMU Health & Medicine
@@
 GMAI-VL(ours) 50.0 60.0 43.3 50.0 53.3 51.3
-5.5 Ablation Study on GMAI-VL-5.5M Dataset
-We further analyzed the effectiveness of the proposed GMAI-VL-5.5M dataset.
-By removing GMAI-VL-5.5M from the training set, we observed a significant performance drop, as shown in Table 2, Table 3, Table 4 and Table 5.
-In the VQA-RAD, SLAKE, and PMC-VQA benchmarks, removing GMAI-VL resulted in an average performance improvement of 8.6%.
-For the OmniMedVQA, GMAI-MMBench, and MMMU Health & Medicine track, the improvements were 8.52%, 6.75%, and 5.3%, respectively.
-These results demonstrate that the GMAI-VL-5.5M dataset provides highly accurate and reliable medical knowledge, particularly in recognizing and understanding multimodal medical data, which significantly boosts model performance.
-This underscores the dataset’s diversity, comprehensiveness, and its ability to effectively complement other datasets in complex medical tasks.
+Figure 2: Example results of our GMAI-VL model. Figure (e) is a failed case.
+The OmniMedVQA [32] benchmark integrates 73 traditional medical imaging datasets, all formatted for visual question answering (VQA).
+Table 5 summarizes the performance of various large vision-language models (LVLMs), including GMAI-VL, across five question types: Modality Recognition, Anatomy Identification, Disease Diagnosis, Lesion Grading, and Other Biological Attributes.
+From the results, we can see that (i) GMAI-VL excels in all tasks, achieving 98.64% in Modality Recognition, 92.95% in Anatomy Identification, and 88.71% in Disease Diagnosis; (ii) it outperforms both open-source and medical-specific models, demonstrating strong abilities in identifying anatomical structures and diagnosing diseases; (iii) in Lesion Grading, it achieves the highest score of 87.21%, and scores 82.95% in Other Biological Attributes, showcasing its versatility; and (iv) with an average accuracy of 88.48%, GMAI-VL surpasses models like HuatuoGPT-Vision-34B and InternVL2-40B, establishing itself as a leading model in multimodal medical image understanding and setting a new benchmark for medical VQA tasks.
+5.3 Comparisons on GMAI-MMBench
+The GMAI-MMBench benchmark is a comprehensive evaluation suite for medical multimodal models, focusing on clinical visual question-answering (VQA) tasks.
+Table 6 presents the performance of various models on the val and test sets across a range of clinical tasks.
+Notably, (i) GMAI-VL excels in tasks such as abnormality recognition (73.78%), biological variation recognition (63.06%), and clinical disease diagnosis (66.67%), demonstrating its strong ability to understand and interpret complex clinical images.
+(ii) Compared to other models, GMAI-VL consistently ranks first or second across most tasks, securing the top position in 16 out of 20 categories.
+(iii) Key tasks such as Attribute Recognition (AR) and Disease Diagnosis (DD) yield scores of 75.26% and 67.14%, respectively, highlighting GMAI-VL’s strength in medical scenario understanding.
+(iv) Overall, GMAI-VL sets a new benchmark in various clinical VQA tasks.
+5.4 Comparisons on MMMU Health & Medicine
+We further assess the performance of our GMAI-VL on the Health & Medicine track of MMMU benchmark, which is a widely recognized standard for evaluating multimodal models.
+The experimental results in Table 8 show the model’s performance across five key categories: Basic Medical Science (BMS), Clinical Medicine (CM), Diagnostics and Laboratory Medicine (DLM), Pharmacy (P), and Public Health (PH).
+The results show that (i) GMAI-VL performs strongly across multiple categories, achieving top scores in DLM (43.3%), P (50.0%), and PH (53.3%), surpassing competitive models such as LLaVA-v1.6 and HuatuoGPT-Vision-7B.
+These results highlight the model’s proficiency in handling complex tasks that require diagnostic reasoning, pharmaceutical knowledge, and public health expertise.
+(ii) In BMS, GMAI-VL scores 50.0%, achieving the best performance and demonstrating its ability to understand medical knowledge.
+(iii) In CM, the model scores 60.0%, remaining competitive with other leading models.
+These results underscore the model’s effectiveness in processing both clinical and foundational medical information.
+(iv) Overall, GMAI-VL achieves an average score of 51.3% across the Health & Medicine track, ranking among the top models and confirming its versatility in specialized medical domains.
+5.5 Ablation Study of the GMAI-VL-5.5M Dataset
+In this section, we evaluate the effectiveness of the proposed GMAI-VL-5.5M dataset.
+We report the results of GMAI-VL (w/o our data) in Tables 4, 5, 6, and 8.
+This model is trained using data excluding the GMAI-VL-5.5M dataset (see Sec.4.2), highlighting that our dataset effectively enhances the performance of large vision-language models.
+The results demonstrate that the GMAI-VL-5.5M dataset provides highly accurate and reliable medical knowledge, especially in recognizing and understanding multimodal medical data.
+This significantly improves model performance, showcasing the dataset’s diversity, comprehensiveness, and its ability to complement other datasets in complex medical tasks.
+5.6 Ablation Study of Training Strategy
+Table 7 presents results at different training stages.
+(i) The proposed stage1+2+3 (Prop) training strategy significantly enhances model performance across multiple benchmarks, outperforming other models and training strategies in all key metrics.
+(ii) By progressively incorporating stages 1, 2, and 3, the model shows consistent improvements on various datasets.
+This highlights the effectiveness of the three-stage training pipeline in improving the model’s ability to handle complex multimodal medical tasks, confirming that a more comprehensive training approach yields superior results.
+5.7 Case Study
+Fig.
+2 shows example results of our GMAI-VL across various medical imaging and diagnostic tasks.
+The model accurately interprets chest X-rays, detects lesions in fundus images, and identifies COVID-19 features in chest CT scans, among others.
+These results highlight the model’s versatility and potential in assisting clinical diagnostics.
+However, as illustrated in the failure case in Fig.
+2 (e), GMAI-VL encounters difficulties in detecting subtle differences, such as distinguishing between mild and moderate severity.
 6 Conclusion
-In this study, we develop GMAI-VL, a large vision-language model, along with GMAI-VL-5.5M, a comprehensive multimodal medical dataset aimed at advancing general medical AI (GMAI).
-GMAI-VL-5.5M, which converts hundreds of medical image analysis datasets into high-quality image-text pairs through annotation-guided data generation, enables GMAI-VL to effectively address a wide range of clinical tasks.
-Experimental results show that GMAI-VL-5.5M significantly enhances GMAI-VL’s performance on diverse clinical tasks, achieving state-of-the-art results across several key benchmarks.
+In this paper, we build GMAI-VL, a large vision-language model, and GMAI-VL-5.5M, a comprehensive multimodal medical dataset designed to advance general medical AI.
+GMAI-VL-5.5M converts hundreds of medical image analysis datasets into high-quality image-text pairs through annotation-guided data generation, enabling GMAI-VL to tackle a wide range of clinical tasks effectively.
+Experimental results show that GMAI-VL-5.5M significantly enhances GMAI-VL’s performance, achieving state-of-the-art results across multiple key benchmark datasets.
+Future work will focus on expanding the dataset with more diverse and challenging medical scenarios, further improving the model’s ability to generalize across different clinical environments and applications.
+{CJK}
+UTF8gbsn
 Supplementary Material
-7 Expanded Data Generation Example
-In this section, we present the full version of Fig.
-2, which provides a detailed example of the data generation process, as shown in Fig.3.
-This expanded illustration includes the complete prompt and the corresponding detailed responses, enabling a clear comparison between data generation methodologies with and without annotation guidance.
-The detailed comparison allows for a more intuitive understanding of the distinctions between both methods.
-Figure 3: The full version of Fig.
-2 in the main text illustrates the complete of data generation pipeline, comparing without-annotation-guided and annotation-guided methods.
-The annotation-guided approach integrates specific annotation information (e.g., <image, modality, label, department, bbox [optional]>) to generate high-quality, accurate descriptions, while the without-annotation-guided approach often results in lower-quality outputs.
-8 Details of GMAI-VL-5.5M
-In Table.
-6, we provide sub-datasets information of the multimodal dataset GMAI-VL-5.5M we have constructed.
-Based on the different data forms introduced in the paper, we have categorized the data into five distinct sub-datasets.
-These include GMAI-MM-Caption-1.7M, GMAI-MM-Instruct-0.9M, GMAI-MM-Percept-1.3M, GMAI-Text-Single-1M, and GMAI-Text-0.6M.
-Each sub-dataset corresponds to specific components: image caption data, free instruction data, visual perception data, text-only, and conversation.
-Furthermore, Fig.
-4 presents a comprehensive distribution of the data across different modalities, original tasks, departments, and clinical task within this multimodal dataset, highlighting its extensive richness and diversity.
-Table 6: Sub-Dataset Details for GMAI-VL-5.5M
+GMAI-VL & GMAI-VL-5.5M: A Large Vision-Language Model and a Comprehensive Multimodal Dataset Towards General Medical AI
+There are four parts in this supplementary material.
+Part 1: Details of GMAI-VL-5.5M
+Part 2: Details of All Training Data for GMAI-VL
+Part 3: Results on the Validation and Test Sets of GMAI-MMBench for Clinical VQA Tasks
+Part 4: Quality Evaluation
+Part 1 - Details of GMAI-VL-5.5M
+Table 9: Details of Sub-Datasets in GMAI-VL-5.5M
 Dataset Sub-Dataset Name Description Size
@@
 GMAI-Text-Multi-0.6M A dataset of multi-turn medical conversations on various topics. 0.6M
-9 Training Data Overview
-In this section, we provide a comprehensive overview of all datasets utilized for training the GMAI-VL model.
-The details include the dataset names, their corresponding categories, the amount of data used for training, and the proportion of training data allocated to each dataset during the three phases of model training.
-Fig.
-5 visualizes the distribution of all our training data, showing the proportion of each category and subcategory.
-Table.
-7 summarizes the datasets employed, highlighting their respective categories and sizes.
-It is important to note that for certain datasets, we performed data cleaning and bilingual translation.
-As a result, the dataset sizes reported here may differ from the official numbers.
 (a) (a) Modality distribution
@@
 (d) (d) Clinical task distribution
-Figure 4: Distribution of GMAI-VL-5.5M across tasks, modalities, departments, and clinical tasks.
-(a) Original Task Distribution: The dataset includes 2D Classification (50.4%), 3D Segmentation (30.3%), 2D Segmentation (12.7%), and 2D Detection (6.6%).
-(b) Modality Distribution: In addition to CT (26.8%) and MR (24.7%), X-ray (12.6%), Pathology (11.2%), and less common modalities like Dermoscopy (3.5%), Microscopy (2.4%), and PET (0.2%) are represented.
-(c) Department Distribution: While Orthopedic Surgery (12.9%) and General Surgery (10.3%) are the top contributors, departments like Endocrinology (1.3%), Infectious Diseases (0.8%), and Urology (0.7%) also provide data.
-(d) Clinical Task Distribution: Besides Disease Diagnosis (40.4%) and Organ Recognition (16.0%), tasks such as Muscle Recognition (3.3%), Nervous Tissue Recognition (1.5%), and Microorganism Recognition (1.2%) are included.
-Note: The distribution statistics shown here pertain only to the multimodal components of GMAI-VL-5.5M.
-Figure 5: Distribution of all our training data.
-The inner ring represents major categories, each depicted in a different color.
-The outer ring corresponds to the subcategories within each major category.
-The size of each segment is proportional to the amount of data, as indicated in the legend, where the data volume for each subcategory is also provided.
+Figure 3: Distribution of GMAI-VL-5.5M across various modalities, departments, and tasks.
+Table 9 provides information on the sub-datasets of the multimodal dataset GMAI-VL-5.5M that we have constructed.
+Based on the different data formats discussed in the paper, we have categorized the data into five distinct sub-datasets: GMAI-MM-Caption-1.7M, GMAI-MM-Instruct-0.9M, GMAI-MM-Percept-1.3M, GMAI-Text-Single-1M, and GMAI-Text-0.6M.
+Each sub-dataset corresponds to specific components, including image caption data, free instruction data, visual perception data, text-only data, and conversation data.
+Additionally, Fig.
+3 presents a comprehensive distribution of the data across different modalities, original tasks, departments, and clinical tasks within the dataset, highlighting its richness and diversity.
+Part 2 - Details of All Training Data for GMAI-VL
+Figure 4: Distribution of the training dataset.
+The inner ring represents major categories, each shown in a distinct color, while the outer ring depicts the corresponding subcategories.
+Segment sizes are proportional to data volume, as indicated in the legend, which also provides the data volume for each subcategory.
 Table 10: List of datasets used in our model. We employ a large collection of image-text data and instruction data for training stage.
@@
 Overall - 15.7M - -
-Figure 6: Diagram of the three-stage training process.
-10 Model Training Settings
-Table 8: Training settings of GMAI-VL’s stage I, stage II, and stage III.
-Settings Stage I Stage II Stage III
-freeze LLM True True False
-freeze MLP False False False
-freeze Vision Encoder True False False
-packing type soft packing soft packing soft packing
-learning rate 1e-3 1e-4 1e-5
-learning rate schedule cosine decay cosine decay cosine decay
-optimizer AdamW AdamW AdamW
-optimizer hyper-parameters $\beta_{1}=0.9,\beta_{2}=0.999$ $\beta_{1}=0.9,\beta_{2}=0.999$ $\beta_{1}=0.9,\beta_{2}=0.999$
-input size 336x336 336x336 336x336
-total batch size 32x8x2 32x4x4 32x4x4
-drop rate 0.0 0.0 0.0
-numerical precision DeepSpeed bf16 DeepSpeed bf16 DeepSpeed bf16
-GPUs for training 32xA100 (80G) 32xA100 (80G) 32xA100 (80G)
-The table. 8 presents the training settings for GMAI-VL across three stages, detailing key hyperparameters. The train diagram as illustrated in Fig. 6.
-  • Stage I (Shallow alignment).
-During this stage, both the large language model (LLM) and the vision encoder are frozen, while the MLP remains trainable.
-The learning rate is set to $1e^{-3}$ with a cosine decay schedule, using AdamW as the optimizer.
-Input size is $336\times 336$, and the total batch size is $32\times 8\times 2$.
-Here, 32 refers to the number of GPUs used, 8 represents the micro-batch size per GPU, and 2 is the number of gradient accumulation steps.
-Notably, we use a soft packing technique, which allows each sample to actually contain multiple sequences—estimated at over 2 per sample.
-  • Stage II (Deep alignment).
-The LLM remains frozen, while the vision encoder and MLP are unfrozen and trainable, allowing for deeper feature alignment.
-The learning rate is lowered to $1e^{-4}$, the total batch size is $32\times 4\times 4$, and other settings remain consistent with Stage I.
-  • Stage III (Instruction tuning).
-In this stage, all components of the model are trainable, allowing for comprehensive fine-tuning.
-The learning rate is further reduced to $1e^{-5}$, the total batch size is $32\times 4\times 4$, while other parameters, including optimizer, remain unchanged across stages.
-Each stage utilizes DeepSpeed for mixed-precision training (bf16) on 32 A100 (80GB) GPUs.
-To accelerate the training rate of the model, we utilize the soft packing method in xtuner***https://github.com/InternLM/xtuner, which packs multiple sequences into each sample to optimize memory usage and efficiency..
-Figure 7: Examples of our model in various medical tasks, including image description, disease diagnosis, diagnostic report generation, free question answering, etc.
-The results demonstrate the model’s versatility across different imaging modalities and tasks, with support for both Chinese and English.
-11 Results
+In this part, we provide a comprehensive overview of all datasets used for training the GMAI-VL model.
+This includes the dataset names, categories, the amount of data used, and the proportion of training data allocated to each dataset during the three phases of model training.
 Fig.
-7 illustrates several examples of our model on various tasks including image description, disease diagnosis, diagnostic report generation, free question answering, etc.
-These examples cover a diverse range of medical imaging modalities and tasks, highlighting the model’s versatility and general capabilities within medical domain.
-Furthermore, GMAI-VL supports both Chinese and English, thereby facilitating bilingual functionality and expanding its applicability in diverse clinical settings.
+4 visualizes the distribution of our training data, illustrating the proportion of each category and subcategory.
+Table below summarizes the datasets employed, along with their respective categories and sizes.
+It is important to note that for some datasets, we performed data cleaning and bilingual translation, which may result in reported dataset sizes differing from the official numbers.
+Part 3 - Results on the Validation and Test Sets of GMAI-MMBench for Clinical VQA Tasks
+For further comparisons, Table 11 provides additional results, including a broader range of models, such as Open-Source LVLMs, Proprietary LVLMs, and Medical Special Models.
+Table 11: Results on the val and test sets of GMAI-MMBench for clinical VQA tasks.
+The full names of the evaluated tasks can be found in Table 5 in literature[16].
+The best model in each category is in red, while the second-best is blue.
+Model Name Overall (val) Overall (test) AR BVR B CR C DD IQG MR M NT OR-A OR-HN OR-P OR-T SG SAR SIR SWR
+Random Guess
+Random 25.70 25.94 38.20 22.73 22.92 22.72 24.06 26.66 27.13 27.00 20.00 24.75 21.37 22.93 22.33 21.18 32.43 24.23 21.39 23.71
+Open-Source LVLMs
+Flamingo v2[4] 25.58 26.34 37.74 21.50 20.62 22.00 22.41 27.29 25.91 27.45 18.00 28.79 25.16 22.13 22.00 22.00 34.61 22.88 20.44 27.43
+VisualGLM-6B[22] 29.58 30.45 40.16 33.92 24.92 25.22 24.21 32.99 29.96 29.53 21.20 37.88 30.32 24.80 13.33 29.88 33.11 19.62 19.16 37.43
+InstructBLIP-7B[20] 31.80 30.95 42.12 26.92 24.92 28.09 21.65 34.58 31.58 29.23 22.40 30.30 28.95 27.47 23.00 24.82 32.88 19.81 21.64 26.57
+Qwen-VL[6] 34.80 36.05 37.05 37.24 35.85 28.98 24.81 43.60 24.70 30.12 19.20 44.44 29.68 31.87 25.00 31.18 30.26 21.54 20.10 26.86
+Yi-VL-6B[74] 34.82 34.31 41.66 39.16 26.62 30.23 31.88 38.01 26.72 24.93 25.20 37.37 29.58 31.20 32.33 30.59 36.71 24.81 23.18 31.43
+ShareGPT4V-7B[15] 36.71 36.70 43.96 37.59 21.54 37.57 18.80 43.26 32.39 27.30 22.80 43.43 29.47 37.33 22.00 31.76 34.98 24.42 25.06 30.00
+LLAVA-V1.5-7B[45] 38.23 37.96 45.45 34.27 30.92 41.32 21.65 44.68 34.01 27.74 23.60 43.43 28.00 42.13 29.00 35.06 33.41 22.12 23.61 29.14
+XComposer2[24] 38.68 39.20 41.89 37.59 33.69 40.79 22.26 45.87 36.44 32.94 27.20 58.59 26.11 36.40 43.67 37.29 32.06 23.46 27.80 32.86
+LLAVA-InternLM-7b[19] 38.71 39.11 36.36 36.54 32.62 38.10 30.68 46.53 34.82 28.19 25.20 48.99 28.11 40.53 33.33 36.00 34.08 26.73 24.12 29.71
+InternVL-Chat-V1.5[18] 38.86 39.73 43.84 44.58 34.00 33.99 31.28 45.59 33.20 38.28 32.40 42.42 31.89 42.80 27.00 36.82 34.76 23.27 24.72 32.57
+InternVL-Chat-V1.2[17] 39.52 40.01 41.66 44.06 27.38 38.46 34.29 46.99 33.60 34.42 21.20 47.98 30.63 42.80 27.67 35.88 35.59 23.85 24.98 28.00
+LLAVA-InternLM2-7b[19] 40.07 40.45 39.82 37.94 30.62 35.24 29.77 48.97 34.01 25.96 20.80 53.03 30.95 42.67 32.00 39.88 32.43 21.73 24.38 38.00
+DeepSeek-VL-7B[49] 41.73 43.43 38.43 47.03 42.31 37.03 26.47 51.11 33.20 31.16 26.00 44.95 36.00 58.13 36.33 47.29 34.91 18.08 25.49 39.43
+MiniCPM-V2[70] 41.79 42.54 40.74 43.01 36.46 37.57 27.82 51.08 28.74 29.08 26.80 47.47 37.05 46.40 25.33 46.59 35.89 22.31 23.44 31.71
+Proprietary LVLMs
+Claude3-Opus[2] 32.37 32.44 1.61 39.51 34.31 31.66 12.63 39.26 28.74 30.86 22.40 37.37 25.79 41.07 29.33 33.18 31.31 21.35 23.87 4.00
+Qwen-VL-Max[6] 41.34 42.16 32.68 44.58 31.38 40.79 10.68 50.53 32.79 44.36 29.20 51.52 41.37 58.00 30.67 41.65 26.95 25.00 24.64 39.14
+GPT-4V[1] 42.50 44.08 29.92 48.95 44.00 37.39 12.93 52.88 32.79 44.21 32.80 63.64 39.89 54.13 37.00 50.59 27.55 23.08 25.75 37.43
+Gemini 1.0[61] 44.38 44.93 42.12 45.10 46.46 37.57 20.45 53.29 35.22 36.94 25.20 51.01 34.74 59.60 34.00 50.00 36.64 23.65 23.87 35.43
+Gemini 1.5[55] 47.42 48.36 43.50 56.12 51.23 47.58 2.26 55.33 38.87 48.07 30.00 76.26 51.05 75.87 46.33 62.24 20.57 27.69 30.54 40.57
+GPT-4o[1] 53.53 53.96 38.32 61.01 57.08 49.02 46.62 61.45 46.56 56.38 34.00 75.25 53.79 69.47 48.67 65.88 33.93 22.88 29.51 39.43
+Medical Special Model
+Med-Flamingo[53] 12.74 11.64 6.67 10.14 9.23 11.27 6.62 13.43 12.15 6.38 8.00 18.18 9.26 18.27 11.00 11.53 12.16 5.19 8.47 11.43
+LLaVA-Med[40] 20.54 19.60 24.51 17.83 17.08 19.86 15.04 19.81 20.24 21.51 13.20 15.15 20.42 23.73 17.67 19.65 21.70 19.81 14.11 20.86
+Qilin-Med-VL-Chat[48] 22.34 22.06 29.57 19.41 16.46 23.79 15.79 24.19 21.86 16.62 7.20 13.64 24.00 14.67 12.67 15.53 26.13 24.42 17.37 25.71
+RadFM[68] 22.95 22.93 27.16 20.63 13.23 19.14 20.45 24.51 23.48 22.85 15.60 16.16 14.32 24.93 17.33 21.53 29.73 17.12 19.59 31.14
+MedDr[29] 41.95 43.69 41.20 50.70 37.85 29.87 28.27 52.53 36.03 31.45 29.60 47.47 33.37 51.33 32.67 44.47 35.14 25.19 25.58 32.29
+Our Model
+GMAI-VL(w/o our data) 54.99 56.23 51.26 61.05 53.79 44.39 44.51 62.60 40.80 57.42 35.20 79.50 61.31 77.81 53.60 69.29 35.39 35.77 29.71 44.86
+GMAI-VL(ours) 61.74 62.43 75.26 59.66 67.24 56.86 54.29 67.14 42.80 79.97 41.60 75.00 60.45 75.48 53.33 58.12 42.09 72.31 37.40 59.14
+Part 4 - Quality Evaluation
+Table 12: Prompt for Evaluating the Quality of Medical Image-Text Dialogues
+You are a medical imaging expert tasked with scoring each medical text-image dialogue according to the following criteria.
+The score range is from 1 to 5, where a lower score indicates higher dialogue quality.
+Each dialogue should be evaluated based on the following criteria, considering language clarity, completeness, medical accuracy, and relevance to medical imaging.
+Below is an example of the dialogue content and related information:
+- Modality: <image modality>(e.g., CT, MRI, X-ray, etc.)
+- Department: <department>(e.g., Ophthalmology, Pulmonary etc.)
+- Label: <label>(e.g., The pre-defined label belong to image)
+- Dialogue Content:
+ - Question: <question>Question content</question>
+ - Answer: <answer>Answer content</answer>
+Please score each dialogue based on the following dimensions:
+1. Medical Information Accuracy:
+ - Ensure that medical terminology, diagnostic processes, and treatment plans align with current medical standards and practices.
+ - Assess whether there are any medical errors or misleading information that could impact diagnostic or treatment decisions.
+ - Example: If an imaging description or medical terminology is incorrect, how does this affect the overall evaluation?
+2. Language Clarity and Fluency:
+ - Evaluate whether the dialogue is concise, clear, and easy to understand.
+ - Assess whether the language is fluent and natural, and if there are any overly complex sentence structures that might hinder information transmission.
+ - Example: Are there any unnecessary words or overly complex structures that slow down the understanding of the content?
+3. Dialogue Completeness:
+ - Evaluate whether the dialogue includes complete medical information, covering the patient’s history, imaging results, diagnostic information, and any necessary details.
+ - Assess whether any essential information is missing, omitted, or incomplete.
+ - Example: If imaging results are not mentioned or the information is incomplete, does this affect the overall evaluation?
+4. Medical Imaging Relevance:
+ - Evaluate whether the analysis of medical imaging information is thorough and accurate.
+ - Ensure that the imaging data is appropriately interpreted, the descriptions are clear and accurate, and they align with the medical diagnosis.
+ - Assess whether the imaging data effectively supports or supplements other medical information.
+ - Example: If the description of the imaging data is inaccurate or incomplete, does it compromise the reliability of the diagnostic outcome?
+Scoring Criteria:
+- 1: The dialogue is concise, accurate, and clear, fully conforming to medical standards.
+The imaging interpretation is appropriate, the information is complete and accurate, and it fully supports the medical diagnosis.
+- 2: The dialogue is generally accurate, the language is clear, and the content is complete.
+There may be slight imperfections in expression or minor inaccuracies in imaging interpretation, but overall quality is high.
+- 3: The dialogue is accurate and complete, but there may be slight room for improvement in certain areas.
+For example, the imaging interpretation could be clearer, or language expression could be slightly more refined.
+Overall quality is good and suitable for training data.
+- 4: The dialogue contains some notable medical errors or unclear language that affects accuracy.
+The imaging interpretation is problematic, or it does not integrate well with other clinical information, leading to incomplete content and a reduction in quality.
+- 5: The dialogue contains errors or is unclear, potentially misleading medical decisions.
+There is a lack of effective imaging interpretation, the information is incomplete, and the language is unclear, which compromises the medical diagnosis.
+The overall quality is poor.
+Output Format:
+Please output the results in JSON format as follows:
+{”score”: ”Quality score”, ”comments”: ”the comments for the quality criteria”,}
+To ensure the quality of the generated image-text pairs, we propose a systematic evaluation process based on a 5-level scoring system.
+The definition of each scoring level is outlined in the prompt (Table 12).
+The primary objective of this process is to assess the quality of image-text pairs from a collection of 219 generated datasets and 42 publicly available open-source datasets.
+To ensure the representativeness and scientific rigor of the evaluation, we adopt a sampling survey methodology.
+Specifically, we randomly select a subset of image-text pairs from each dataset for quality assessment, from which we infer the overall quality of the entire dataset.
+This sampling survey strategy enables us to efficiently and scientifically evaluate the quality of the image-text pairs, maximizing the reliability and accuracy of the results while significantly reducing the manual review workload.
+The evaluation process is conducted in multiple stages:
+  1. 1.
+    Initial Quality Assessment: From each dataset, 30 random image-text pairs are selected for evaluation.
+The first step involves an initial quality screening using a multimodal large language model (such as GPT-4o) to automatically assess the quality of the pairs.
+The model utilizes a predefined prompt, as illustrated in Table 12, to assign preliminary quality scores based on the defined criteria (accuracy of medical information, clarity and fluency of language, completeness of the dialogue, and relevance to medical imaging).
+  2. 2.
+    Expert Review: Following the model’s initial assessment, the 30 image-text pairs are then reviewed by five medical imaging experts.
+These experts validate the preliminary scores and provide their independent judgments on the quality of the pairs, ensuring that the evaluation aligns with current medical standards and practices.
+  3. 3.
+    Final Scoring: After expert review, the final score for each dataset is determined based on the consensus of the expert evaluations.
+If the average score of the 30 image-text pairs from a dataset is less than or equal to 3 (on a scale of 1 to 5), the dataset is classified as having satisfactory quality.
+Conclusion: After rigorous evaluation, 95% (247/261) of the datasets met the predefined quality standards and were deemed acceptable.
+Our analysis showed that most datasets generated by GPT using the given labels exhibited high quality, with only a few requiring prompt optimization and additional information to meet the desired standards.
+For low-quality datasets, we addressed the issues by refining the prompt design, improving label specificity, and incorporating additional contextual information to ensure the generated quality met the requirements.
+In contrast, 14 open-release multimodal medical datasets, such as, PMC-OA[43], quilt-1m[33], often extracted from research papers or manually converted, exhibited various issues such as incomplete dialogues, misalignment between images and text, overly brief responses, inclusion of non-medical images, and poor image clarity.
+Due to these limitations, only a portion of the open-source datasets met the required quality standards.
+This systematic approach ensures that only high-quality image-text pairs are included for further use in medical imaging research, with clear documentation of the evaluation process and criteria.
+{CJK}
+UTF8gbsn